### v1

Abstract
With the rapid development of machine learning, numerical discretization methods based on deep neural networks have been widely used in many fields, especially in solving high-dimensional problems where traditional methods face bottlenecks. However, for low-dimensional problems, existing machine learning methods are not as accurate and efficient as traditional methods. In this paper, we propose a subspace method based on neural networks for eigenvalue problems with high accuracy and low cost. Our basic idea is to use neural network based basis functions to span a subspace, then calculate the parameters of the neural network based basis functions through appropriate training, and finally calculate the Galerkin projection of the eigenvalue problem onto the subspace and obtain an approximate solution. In addition, we reduce the dimension of the subspace by applying some dimensionality reduction technique, which can improve the accuracy of the approximate solution in further. Numerical experiments show that we can obtain approximate eigenvalues with accuracy of $10^{-11}$ but with less then 400 epochs, which is significantly superior to other existing neural network based methods.
keywords:
subspace, basis function, eigenvalue problem, multi-eigenpairs, neural networks, high-accuracy
1 Introduction
Eigenvalue problems play an important role in various fields of mathematics, physics, and chemistry, such as material mechanics, fluid mechanics, structural mechanics, and quantum mechanics. Accordingly, the study of efficient numerical discretization methods to tackle these problems is an active field in scientific computing. Over recent decades, many numerical methods have been developed to discretize eigenvalue problems with great success, including the finite difference method [26], the finite volume method [15, 31], the finite element method [6, 9, 41], and the spectral method [20].
With the rapid development of machine learning, many numerical discretization methods for partial differential equations (PDEs) based on deep neural networks have received increasing attention, such as the physical-information neural networks (PINN) [35], the deep Galerkin method (DGM) [39], the deep Ritz method (DRM) [50], the deep Nitsche method (DNM) [32], and the weak adversarial networks (WAN) [52]. By these methods, the solution of the partial differential equation is represented by a neural network directly and the optimal parameters of the neural network is obtained by minimizing a loss function. The main difference between these methods lies in the choice of the loss function. For instance, PINN and DGM use the $L^{2}$ norm of the residual in the strong form as the loss function, while DRM and WAN use the energy functional and the operator norm in the weak form as the loss function, respectively. In contrast, DNM uses a loss function based on Nitsche’s energy formulation to handle essential boundary conditions.
Due to the powerful approximation capabilities of neural networks, the aforementioned methods are not only used to deal with source problems but also to handle eigenvalue problems. For instance, in [22], a wave function ansatz based on deep neural networks is introduced to solve the many-electron Schrödinger equation. In [23, 24], various wave function ansatzes in the form of a Slater determinant are also introduced for the Schrödinger equation. In [28], a neural network based multistate solver is proposed for finding the first $k$ eigenpairs of the Schrödinger equation simultaneously. In [21], a diffusion Monte Carlo-like method is proposed for handling high-dimensional eigenvalue problems. In [27], a semigroup method is provided for solving high-dimensional elliptic partial differential equations using neural networks and is extended to a primal-dual algorithm for solving eigenvalue problems. In [29], a type of tensor neural network based on the DECOMP/PARAFAC (CP) tensor decomposition is proposed, along with an efficient numerical integration scheme for tensor-type functions. Theoretical analyses and applications of tensor neural networks are provided in [44] and [43, 45], respectively. For example, tensor neural networks have been used to solve the the first eigenvalue and the corresponding eigenfunction of Schrödinger equation in [43] and the first $k$ eigenpairs of high-dimensional eigenvalue problems in [45]. Additionally, in [51], a machine learning method is proposed for computing eigenvalues and eigenfunctions of the Schrödinger equation, which employs trial functions satisfying the homogeneous boundary conditions to avoid the penalty of the boundary, and an explicit convergence rate of the generalization error also is provided in the article.
While deep neural network based methods have made strides in solving high-dimensional problems where traditional methods fail, they face challenges in solving low-dimensional problems due to the lack of algorithms to train neural networks efficiently. As a result, they lag behind traditional methods in accuracy and efficiency, which prevent most deep neural network based methods from being not only usable but also efficient.
In recent years, some emerging methods based on shallow neural networks have received increasing attention, such as the extreme learning machine (ELM) [25] and the random feature method (RFM) [10, 13]. In these methods, all parameters of the neural network except for the output layer are randomly generated and fixed, which simplifies the optimization problem into a linear system. By using the least-squares method to solve the linear system, the expensive training process can be carried out at a low cost with even higher accuracy. These methods and their variants [17, 38, 40] have been widely used for solving partial differential equations [10, 12, 30, 46]. However, as so far, there is little application of these methods for solving eigenvalue problems.
From another perspective, the methods above fundamentally offer a method to construct basis functions using neural networks. However, as the basis functions constructed by neural networks are sensitive to the choice of hyperparameters, improper choices can significantly impact the accuracy of the approximate solution. Randomly generating these hyperparameters is proved to be inefficient for some problems, which means that the search for more reasonable ways to select and update hyperparameters is a worthwhile pursuit.
To address the impact of hyperparameter selection on solutions, a modified batch intrinsic plasticity method for pre-training the randomly generated hyperparameters in ELM is proposed in [18]. A differential evolution algorithm is employed to compute the optimal or near-optimal parameter $R_{m}$, where the parameters of the neural network are set to random values generated in interval $[-R_{m},R_{m}]$ in [19]. Additionally, a better way to randomly select hyperparameters based on some geometric explanations is provided in [53]. There are also some methods that couple the updating of hyperparameters with the solution of the linear system. An alternative iteration between gradient descent to update the hyperparameters and least squares to solve the linear system is adopted in [2, 14].
Recently, a subspace method based on neural networks (SNN) for partial differential equations in strong form has been proposed in [49], which provides a deterministic way to obtain the appropriate hyperparameters. The main idea of the method is to use neural network based functions as basis functions to span a subspace, then obtain the parameters of the neural network based functions by minimizing a loss function and subsequently find the approximate solution in the subspace. Unlike ELM-like methods and random feature methods, parameters of the neural network in SNN are not randomly generated; instead, they are obtained by minimizing some loss function. In comparison to classical deep neural network based discretization methods, the training process in the SNN focuses on finding better basis functions, which requires relatively lower accuracy in the training process and often needs significantly fewer training epochs. Consequently, SNN can achieve high accuracy with low cost in training for solving partial differential equations. Furthermore, a weak form version of SNN has also been proposed in [33].
From another perspective, the training process of SNN essentially offers a deterministic approach to generate a basis based on neural networks. The solving process of SNN, which calculate the Galerkin projection of the original problem onto the subspace spanned by the basis functions obtained from the training process, can be viewed as a special case of finite dimensional approximation commonly employed in traditional methods. The difference lies in the selection of basis functions: traditional methods like the finite element method employs piecewise polynomials as basis functions, spectral method employs orthogonal polynomials as basis functions, whereas SNN employs appropriately trained neural network based functions as basis functions.
In this paper, motivated by the work in [49], we first propose a subspace method based on neural networks for eigenvalue problems and then an improved version. In the improved version, we add a dimensionality reduction step after obtaining the subspace spanned by the neural network based basis functions. Our method mainly consist of four steps. First, we specify neural networks architecture and appropriately modify basis functions as needed for the specific problem. Next, we obtain the parameters of the basis functions by minimizing some loss function. Then, we reduce the dimension of the subspace spanned by the basis functions by some dimensionality reduction technique. Finally, we calculate the Galerkin projection of the original eigenvalue problem onto the reduced subspace and obtain an approximate solution.
We apply our method to solution of some typical eigenvalue problems. In those examples, our algorithm can achieve very high accuracy with few training epochs, which demonstrates the advantages of our new algorithm in both computational cost and accuracy.
The subsequent parts of our paper are organized as follows. In Section 2, we provide some preliminary knowledge, including some notation and classical results for a linear elliptic eigenvalue problem, and an briefly introduction of the subspace method based on neural networks for some source problems. In Sections 3, we introduce the general framework of the subspace method based on the neural networks for eigenvalue problems and the details for each step. In Section 4, we introduce some dimensionality reduction technique to reduce the dimension of the subspace spanned by the neural network based functions, and obtain the improved version of SNN for eigenvalue problems. In Section 5, we provide some numerical results to show the efficiency of our new method. Finally, we conclude with a brief conclusion.
2 Preliminaries
Before introducing our method, we first give some background knowledge, including some basic results of a elliptic eigenvalue problem and some briefly introduction to the subspace method based on neural networks for partial differential equations proposed in [49].
2.1 A linear elliptic eigenvalue problem and its finite dimensional discretization
[…]
Consider the following elliptic eigenvalue problem: Find $(\lambda,u)\in\mathbb{R}\times V$ such that $b(u,u)=1$ and
$a(u,v)=\lambda b(u,v),\quad\forall v\in V.$ (1)
[…]
$a(u,v)\lesssim\|u\|\ \|v\|,\forall u,v\in V,$
$\|v\|^{2}\lesssim a(v,v),\forall v\in V\ \text{and}\ 0<b(v,v),\forall v\in V,v\neq 0,$
In further, we define the norms $\|\cdot\|_{a}$ and $\|\cdot\|_{b}$ based on the bilinear forms $a(\cdot,\cdot)$ and $b(\cdot,\cdot)$ as follows:
$\|\cdot\|_{a}=\sqrt{a(\cdot,\cdot)},\|\cdot\|_{b}=\sqrt{b(\cdot,\cdot)}.$
[…]
It is well known that the eigenvalue problem (1) has an eigenvalue sequence $\{\lambda_{k}\}$ ([6])
$0<\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{k}\leq\cdots,\lim_{k\to\infty}\lambda_{k}=\infty,$
and corresponding eigenfunctions
$u_{1},u_{2},\cdots,u_{k},\cdots,$
which satisfy
$b(u_{i},u_{j})=\delta_{ij},i,j=1,2,\cdots,k,\cdots.$
In the sequence $\{\lambda_{k}\}$, $\lambda_{k}$ are repeated according to their geometric multiplicity.
Let $V_{h}$ be a finite dimensional subspace of $V$. The finite dimensional discretization of (1) in $V_{h}$ is as follows: Find $(\lambda,u_{h})\in\mathbb{R}\times V_{h}$ such that $b(u_{h},u_{h})=1$ and
$a(u_{h},v_{h})=\lambda b(u_{h},v_{h}),\quad\forall v_{h}\in V_{h}.$ (2)
Suppose $V_{h}=\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$, where $\varphi_{1},\cdots,\varphi_{M}$ are basis functions of the finite dimensional subspace $V_{h}$. Substituting $u_{h}=\sum_{j=1}^{M}w^{(j)}\varphi_{j}\ \text{and choosing}\ v_{h}\ \text{to be}\ \varphi_{i}$ into (2), we then have:
$\sum\limits_{j=1}^{M}w^{(j)}a(\varphi_{j},\varphi_{i})=\lambda\sum\limits_{j=1}^{M}w^{(j)}b(\varphi_{j},\varphi_{i}),i=1,\cdots,M.$
Denote $A$ the stiffness matrix and $B$ the mass matrix, that is, $A=(a(\varphi_{i},\varphi_{j}))_{ij}\in\mathbb{R}^{M\times M},B=(b(\varphi_{i},\varphi_{j}))_{ij}\in\mathbb{R}^{M\times M}$. The eigenvalue problem (1) can be formulated as the following matrix eigenvalue problem:
$Aw=\lambda Bw,$ (3)
where $w=(w^{(1)},\cdots,w^{(M)})^{T}\in\mathbb{R}^{M}.$
By solving the matrix eigenvalue problem (3), an approximate solution to the eigenvalue problem (1) can be obtained.
2.2 Subspace method based on neural networks
In this subsection, we will give a brief review for the subspace method based on neural networks proposed in [49] for the partial differential equations, which has been proven to be able to achieve high accuracy with low cost.
The neural networks architecture used in [49] is comprised of four parts: an input layer, several hidden layers, a subspace layer and an output layer. For simplicity, we take the case of single output as an example to show the mathematical expression of the neural network. The expression of multiple outputs can be obtained similarly.
Denote $L$ the number of hidden layers, $n_{0}=d$ the number of neurons in the input layer, $n_{1},\cdots,n_{L}$ the number of neurons in each hidden layer, $n_{L+1}=M$ the number of neurons in the subspace layer, and $n_{L+2}=1$ the number of neurons in the output layer. The neural network used in [49] can be expressed as follows:
[…]
Here, $W_{l}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ and $\bm{b}_{l}\in\mathbb{R}^{n_{l}}$ are the weights and biases of the corresponding layer, respectively, $\bm{x}\in\mathbb{R}^{d}$ is the input, $\varphi=(\varphi_{1},\cdots,\varphi_{M})^{T}$ and $w=(w^{(1)},\cdots,w^{(M)})^{T}$, $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ is an activate function which can be extended to $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ as follows:
$\sigma(\bm{x})=\begin{bmatrix}\sigma(x_{1})\\ \vdots\\ \sigma(x_{n})\end{bmatrix}\ \text{for any}\ \bm{x}=(x_{1},\cdots,x_{n})^{T}\ \in\mathbb{R}^{n}.$
Denote $\theta=\{W_{1},\cdots,W_{L+1},\bm{b}_{1},\cdots,\bm{b}_{L+1}\}\in\Theta:=\mathbb{R}^{n_{1}\times n_{0}}\times\cdots\times\mathbb{R}^{n_{L+1}\times n_{L}}\times\mathbb{R}^{n_{1}}\times\cdots\times\mathbb{R}^{n_{L+1}}$ the parameters in the neural network, then $u(\bm{x};\theta,w)$ is the output of $\bm{x}$ with respect to parameters $\theta$ and $w$.
Next, we will review the general framework of the subspace method based on neural networks for the partial differential equations proposed in [49].
Firstly, the neural networks architecture introduced above is adopted, and $\theta$ and $w$ are initialized. Next, fix the coefficients $w$, and update the subspace basis functions, i.e., update parameters $\theta$ by minimizing a loss function, aiming to obtain basis functions that are better than the initial ones through training. After training, calculate the Galerkin projection of the original problem onto the subspace spanned by trained basis functions $\varphi_{1},\cdots,\varphi_{M}$ and obtain an approximate solution to the partial differential equation. The above process is summarized as a general framework for solving the partial differential equations in [49]. We also list it here for the completeness, see Algorithm (1) for the details.
Algorithm 1 General framework of SNN for PDEs
  1. Initialize the neural networks architecture (4), generate randomly $\theta$, and give $w$.
  2. Update parameters $\theta$ by minimizing a loss function, and obtain the updated basis functions $\varphi_{1},\cdots,\varphi_{M}$.
  3. Calculate the Galerkin projection of the PDE onto the subspace $\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$ and obtain an approximate solution.
3 Subspace method based on neural networks for eigenvalue problems
We consider the following typical eigenvalue problem:
$\begin{cases}Lu=\lambda u&\text{in }\Omega,\\ u=0&\text{on }\partial\Omega,\end{cases}$
where the differential operator $L:=-\nabla\cdot(\alpha\nabla)+\beta$ and $\alpha=(\alpha_{ij})_{ij}$ is symmetric positive definite with $\alpha_{ij}\in W^{1,\infty}(\Omega)(i,j=1,\cdots,d)$, and $0\leq\beta\in L^{\infty}(\Omega).$
The associate weak form of the eigenvalue problem above is as follows: Find $(\lambda,u)\in\mathbb{R}\times H_{0}^{1}(\Omega)$ such that $b(u,u)=1$ and
$a(u,v)=\lambda b(u,v),\quad\forall v\in H_{0}^{1}(\Omega),$ (5)
which is nothing but (1) with $a(u,v)=(\alpha\nabla u,\nabla v)+(\beta u,v)$ and $b(u,v)=(u,v)$. Here, $a(\cdot,\cdot),b(\cdot,\cdot)$ are two positive definite symmetric bilinear forms on $H_{0}^{1}(\Omega)\times H_{0}^{1}(\Omega)$.
Remark 1
It is worth mentioning that our method is also valid for a more general bilinear form $a(\cdot,\cdot)$ that
$\|v\|_{1}^{2}-C\|v\|_{0}^{2}\lesssim a(v,v),\quad\forall v\in H_{0}^{1}(\Omega)$
holds for some constant $C$. (see, e.g., Remark 2.9 in [16])
3.1 A general framework of SNN for eigenvalue problems
In this subsection, we extend the subspace method based on neural networks for PDEs, which is proposed in [49] and introduced in Section 2, to eigenvalue problems.
A neural networks architecture similar to the one introduced in the previous section is adopted. The only difference is that the neural network in Section 2 is used to approximate the solution to the boundary value problem, here the neural network is used to approximate the eigenfunctions corresponding to the first $k$ eigenvalues of some partial differential operator. Therefore, the dimension of the output layer is $k$. Thus, the output of the neural network can be presented as follows,
$\bm{u}(x;\theta,\bm{w})=\begin{pmatrix}u_{1}(x;\theta,w_{1})\\ \cdots\\ u_{k}(x;\theta,w_{k})\end{pmatrix}=\begin{pmatrix}\varphi\cdot w_{1}\\ \cdots\\ \varphi\cdot w_{k}\end{pmatrix}=\begin{pmatrix}\sum\limits_{j=1}^{M}w_{1}^{(j)}\varphi_{j}(x,\theta)\\ \cdots\\ \sum\limits_{j=1}^{M}w_{k}^{(j)}\varphi_{j}(x,\theta)\end{pmatrix}$
where $\bm{w}=(w_{1},\cdots,w_{k})$, $w_{s}=(w_{s}^{(1)},\cdots,w_{s}^{(M)})^{T}$ denote the coefficients corresponding to the $s$-th output, and $u_{s}(\bm{x};\theta,w_{s})$ denotes the $s$-th output, $s=1,\cdots,k$.
Following a framework similar to Algorithm 1, we initialize $\theta$ and $w_{1},\cdots,w_{k}$ firstly. Then, to handle the boundary condition, appropriate modifications to the basis functions are made according to the specific problem requirements. Next, we fix the coefficients $w_{1},\cdots,w_{k}$ and update parameters $\theta$ by minimizing some loss function, from which we obtain the updated basis functions $\varphi_{1},\cdots,\varphi_{M}$. Finally, we calculate the Galerkin projection of the original eigenvalue problem onto the subspace spanned by basis functions $\varphi_{1},\cdots,\varphi_{M}$ and obtain an approximate solution to the eigenvalue problem (5). The above process is summarized as Algorithm 2.
Algorithm 2 General framework of SNN for eigenvalue problems
  1. Initialize the neural networks architecture (4), generate randomly $\theta$, and give $w_{1}$$\cdots$,$w_{k}$, then modify the neural network basis functions according to the boundary conditions.
  2. Update parameters $\theta$ by minimizing a loss function, and obtain the updated basis functions $\varphi_{1},\cdots,\varphi_{M}$.
  3. Calculate the Galerkin projection of the eigenvalue problem (5) onto the subspace $\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$ and obtain an approximate solution.
In the following subsections, we will follow the process outlined in the Algorithm 2 to provide a step-by-step explanation.
3.2 Initializing and modifying basis functions of the subspace
3.2.1 Initializing the basis functions
First, we introduce the initialization of the neural network parameters $\theta$ and the coefficients $w_{1},\cdots,w_{k}$. There are various ways to initialize the parameters $\theta$ for initializing the neural network. In this paper, we adopt the default initialization strategy provided by PyTorch for neural network parameters $\theta$. The coefficients $w_{1},\cdots,w_{k}$ are given and will remain fixed during the training process. While there are various methods for assigning the coefficients $w_{1},\cdots,w_{k}$, numerical experiments provided in [49] indicate that the different choices of $w_{1},\cdots,w_{k}$ have no obvious impact on the results. Therefore, here, we do not pay much attention on the choice of $w_{1},\cdots,w_{k}$. We just give some simple choices. That is, for a single output, we choose $w=(1,\cdots,1)^{T}\in\mathbb{R}^{M}$. For multiple outputs, a random approach is chosen, that is, $w_{1},\cdots,w_{k}\in\mathbb{R}^{M}$ are assigned random values generated from a uniform distribution defined over the interval $[-1,1]$.
3.2.2 Modifying the basis functions
For a partial differential equation, some boundary conditions are usually required. Since the basis functions constructed by some neural networks are usually not local, it is not so easy to deal with the boundary condition by using the same way as traditional methods. Here, we choose to modify the basis functions to make them satisfy the boundary condition. For example, for the case of zero boundary condition with $\Omega$ being bounded, we can modify the basis function by the following way. We multiply $\varphi_{i}$ by a function $f$ such that
$\bar{\varphi}_{i}=\varphi_{i}f,\ \bar{\varphi}_{i}(\bm{x})=0,\ \forall\bm{x}\in\partial\Omega,\ \forall i=1,\cdots,M.$
For example, for the case of the 2D unit square domain, we can choose $f(\bm{x})=x_{1}x_{2}(1-x_{1})(1-x_{2})$.
If $\Omega$ is unbounded, it is necessary to modify the basis functions to ensure that the modified functions are $L^{2}$ integrable. We can achieve this by multiplying $\varphi_{i}$ by a function $f$ such that
[…]
Since the Hermite-Gauss quadrature scheme are used to numerically approximate the integrals on unbounded domain in this paper, we can choose $f(\bm{x})=\exp(-\frac{1}{2}\bm{x}^{T}\bm{x})$.
For simplicity, in the remaining parts of this paper, $\varphi_{i}$ will refer to the modified basis function $\bar{\varphi}_{i}$ as described above.
3.3 Updating the basis functions
In this subsection, we will introduce a specific approach to update basis functions by minimizing some loss function $\mathcal{L}$.
Denote $\bm{w}=(w_{1},\cdots,w_{k})$. For given coefficients $\bm{w}$, parameters $\theta$ can be updated by solving the following minimization problem:
$\min_{\theta\in\Theta}\mathcal{L}(\theta,\bm{w}).$ (6)
The gradient descent method is adopted to solve minimization problem (6). That is, we update the parameter by
$\theta^{(l+1)}=\theta^{(l)}-\eta\nabla_{\theta}\mathcal{L}(\theta^{(l)},\bm{w}),$
where $\theta^{(l)}$ is the value of $\theta$ in $l$-th epoch and $\eta$ is the learning rate.
We now introduce our choices for the loss function and the termination criterion used in the above process.
Before introducing the loss function we used, we first introduce some properties of the eigenvalue problem (5) to give us an idea on how to choose the loss function.
Denote $V_{k}=\text{span}\{v_{1},\cdots,v_{k}\}$ an arbitrary $k$-dimensional subspace of $V$ and $U_{k}=\text{span}\{u_{1},\cdots,u_{k}\}$ the eigensubspace corresponding to the first $k$ eigenvalues.
According to the the minimum-maximum principle [6], the sum of the first $k$ eigenvalues and eigenfunctions corresponding to the first $k$ eigenvalues satisfy the following equations,
$\sum_{i=1}^{k}\lambda_{i}=\min_{V_{k}=\text{span}\{v_{1},\cdots,v_{k}\}\atop v_{k}\in V,dim(V_{k})=k}\text{Trace}\left(\mathcal{B}^{-1}(v_{1},\cdots,v_{k})\mathcal{A}(v_{1},\dots,v_{k})\right),$ (7)
$\left(u_{1},\cdots,u_{k}\right)=\mathop{\arg\min}_{V_{k}=\text{span}\{v_{1},\cdots,v_{k}\}\atop v_{k}\in V,dim(V_{k})=k}\text{Trace}\left(\mathcal{B}^{-1}(v_{1},\cdots,v_{k})\mathcal{A}(v_{1},\dots,v_{k})\right).$ (8)
Here, $\mathcal{A}$ is the stiffness matrix and $\mathcal{B}$ is the mass matrix, that is, $\mathcal{A}(v_{1},\dots,v_{k})=(a(v_{i},v_{j}))_{ij}\in\mathbb{R}^{k\times k}$, $\mathcal{B}(v_{1},\dots,v_{k})=(b(v_{i},v_{j}))_{ij}\in\mathbb{R}^{k\times k}$.
Motivated by (7) and (8), we then construct the following loss function for finding first $k$ eigenpairs (see also in [45]):
$\mathcal{L}(\theta,\bm{w})=\text{Trace}(\mathcal{B}^{-1}(v_{1}(x;\theta,w_{1}),\cdots,v_{k}(x;\theta,w_{k}))\mathcal{A}(v_{1}(x;\theta,w_{1}),\cdots,v_{k}(x;\theta,w_{k}))),$
where $v_{i}(x;\theta,w_{i})=\sum\limits_{j=1}^{M}w_{i}^{(j)}\varphi_{j}(x,\theta)$ is an approximation of the eigenvector corresponding to the $k$-th eigenvalue, $i=1,\cdots,k.$
As for the choice of termination criteria, we notice that the goal of training is to obtain improved basis functions compared to the initial ones, there is no need to train until the loss is small enough, e.g. find the $\theta^{\ast}$ s.t. $\mathcal{L}(\theta^{\ast},\bm{w})-\min\limits_{\theta\in\Theta}\mathcal{L}(\theta,\bm{w})\leq\epsilon$. Instead, in [49], the following termination criterion is adopted:
$\frac{\mathcal{L}(\theta,\bm{w})}{\mathcal{L}_{0}(\theta,\bm{w})}\leq\epsilon\ \text{or}\ N_{epoch}\geq N_{max},$
where $\mathcal{L}_{0}(\theta,\bm{w})$ is the initial loss, $\epsilon$ is the tolerance, $N_{epoch}$ and $N_{max}$ are the current and maximum number of training epochs, respectively. However, since the minimum value of the loss function we used in this paper is not 0, this criterion may not be suitable for determining when to stop training. As an alternative, we propose a new termination criterion.
We note that for an iterative method, the error between two adjacent iterations is usually used to determine when to terminate the iteration. That is, the iteration is terminated when the error between two adjacent iterations is less than some given tolerance $\epsilon$. Note that the loss usually decreases in an oscillatory manner during the training, hence using the change of the loss function during multiple adjacent epochs as an indicator may be better than using the change during two adjacent epochs. Therefore, we choose to use the change of the loss during multiple adjacent epochs as termination criterion. In this paper, we adopt the change of the loss during 10 adjacent epochs as an indicator. Any other reasonable choices are also acceptable.
Let $l_{s}=\mathcal{L}_{s}(\theta,\bm{w})$ be the loss in the $s$-th epoch, $\bar{l}_{s}=\sum\limits_{i=1}^{10}|l_{s-i}$ $-l_{s+1-i}|$ be the change of the loss function during 10 adjacent epochs. The the following termination criterion is adopted:
$|\frac{\bar{l}_{s}}{l_{s}}|\leq\epsilon,\ \text{or}\ N_{epoch}\geq N_{max},$ (9)
where $\epsilon$ is a given tolerance, $N_{epoch}$ and $N_{max}$ are the current and maximum training epochs, respectively. Under this criterion, the termination of training means that the relative change of the loss function during 10 adjacent epochs has been relatively small, which indicates that the benefits of continued training may be small. Since there is no need to train until the loss is small enough, a relatively big $\epsilon$ and a relatively small $N_{max}$ can be chosen.
3.4 Galerkin projection of the original eigenvalue problem onto the subspace
In this subsection, we will introduce the details on how to obtain the finite dimensional approximation of the eigenvalue problem (3) in the subspace $\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$. From (3) in Section 2, we know that finding the Galerkin approximation of the eigenvalue problem (5) in subspace $,\text{span}\{\varphi_{1},\cdots$$,\varphi_{M}\}$ can be formulated as the following matrix eigenvalue problem:
$Aw=\lambda Bw,$ (10)
where $A=(a(\varphi_{i},\varphi_{j}))_{ij},B=(b(\varphi_{i},\varphi_{j}))_{ij}\in\mathbb{R}^{M\times M},w=(w^{(1)},\cdots,w^{(M)})^{T}\in\mathbb{R}^{M}.$
To calculate the stiffness matrix and mass matrix, we have to do some numerical integration. For simplicity, we take $a(u,v)=(\nabla u,\nabla v),b(u,v)=(u,v)$ as an example to show how to compute each element of the stiffness matrix and mass matrix by numerical integration. They are calculated as follows:
$(\varphi_{i},\varphi_{j})\approx\sum\limits_{l=1}^{n}\rho_{l}\varphi_{i}(\bm{x}^{(l)})\varphi_{j}(\bm{x}^{(l)}),$ (11)
$(\nabla\varphi_{i},\nabla\varphi_{j})\approx\sum\limits_{l=1}^{n}\rho_{l}\nabla\varphi_{i}(\bm{x}^{(l)})\nabla\varphi_{j}(\bm{x}^{(l)}).$ (12)
Here, $\{\bm{x}^{(l)}\}_{l=1}^{n}$ are the numerical integration points, $\{\rho_{l}\}_{l=1}^{n}$ are the weights for each numerical integration point and $n$ is the number of numerical integration points.
The first $k$ eigenpairs $(\lambda_{h,1},w_{1}),\cdots,(\lambda_{h,k},w_{k})$ can be obtained by solving the matrix eigenvalue problem (10) using some eigensolver [8, 37], then the approximation of the first $k$ eigenpairs $(\lambda_{h,1},u_{h,1}),\cdots,$ $(\lambda_{h,k},u_{h,k})$ are obtained by setting $u_{h,i}=\sum\limits_{j=1}^{M}w_{i}^{(j)}\varphi_{j}(x,\theta),i=1,\cdots,k.$
3.5 Limitations
Unfortunately, some simple tests show that simply applying Algorithm 2 to solution of some eigenvalue problems can not obtained results as accurate as solution of some source problems. In fact, we have applied the Algorithm 2 to the 2D Laplace eigenvalue problem on a square domain $[0,1]^{2}$ as introduced in Section 6, aiming to solve the first 15 eigenpairs. The neural networks architecture is constructed by a fully connected neural network with 3 hidden layers and a subspace layer with each hidden layer containing 100 neurons and subspace layer containing $M$ neurons. We choose $\epsilon=10^{-3}$ and $N_{max}=5000$ in the termination criterion and adopt Legendre-Gauss integration scheme with 32 integration points in each dimension. For solving the matrix eigenvalue problem (10), we adopt the solver $scipy.linalg.eig$ in Python. Numerical experiments for different $M$ have been carried and the results are presented in Table 1. We denote $err_{\lambda,l}$ the relative error the $l$-th eigenvalue, which is defined by $err_{\lambda,l}:=\frac{|\lambda_{l}-\lambda_{h,l}|}{|\lambda_{l}|}$, where $\lambda_{l}$ is the reference eigenvalue.
From Table 1 we can see that even when the dimension of the subspace $M$ is small, our method can still obtain approximate eigenvalues with high accuracy. However, as the increase of $M$, the accuracy of eigenvalues is not improved but rapidly deteriorates, contrary to our expectations.
Table 1: The relative errors of the first 15 eigenvalues obtained by Algorithm 2 with different $M$ for the 2D Laplace eigenvalue problem on a square domain $[0,1]^{2}$.
Dimension 100 200 300 400 500 600
$err_{\lambda,1}$ 3.901e-12 1.315e-11 3.477e-04 2.236e-04 4.817e-10 3.047e-04
$err_{\lambda,2}$ 1.586e-11 2.618e-09 9.126e-05 5.287e-05 1.593e-08 3.785e-10
$err_{\lambda,3}$ 5.342e-11 4.210e-03 1.255e-03 1.851e-02 1.476e-01 5.980e-02
$err_{\lambda,4}$ 1.197e-09 6.017e-04 1.471e-06 3.563e-04 1.745e-04 5.853e-04
$err_{\lambda,5}$ 9.387e-10 5.089e-12 8.218e-06 4.899e-03 1.009e-03 8.677e-03
$err_{\lambda,6}$ 5.093e-09 7.511e-11 9.078e-03 1.221e-02 1.116e-01 1.162e-02
$err_{\lambda,7}$ 5.447e-09 3.872e-10 8.271e-06 7.278e-02 3.229e-04 3.899e-03
$err_{\lambda,8}$ 8.941e-09 3.102e-01 6.478e-03 2.485e-01 2.230e-03 4.813e-03
$err_{\lambda,9}$ 3.880e-08 5.917e-02 3.574e-04 9.198e-03 4.877e-03 6.637e-03
$err_{\lambda,10}$ 8.691e-08 1.811e-01 4.090e-02 1.514e-02 2.628e-02 2.852e-02
$err_{\lambda,11}$ 6.166e-08 3.851e-01 2.117e-03 4.508e-03 2.238e-02 9.721e-02
$err_{\lambda,12}$ 2.687e-07 3.014e-01 1.180e-03 1.608e-03 4.303e-02 1.707e-02
$err_{\lambda,13}$ 4.700e-07 4.493e-01 5.860e-02 9.130e-03 5.374e-02 4.256e-02
$err_{\lambda,14}$ 3.048e-07 2.871e-01 5.064e-03 9.433e-03 6.394e-03 3.686e-03
$err_{\lambda,15}$ 7.798e-07 4.797e-01 8.226e-02 9.797e-03 4.927e-02 1.188e-02
After some analysis and numerical verification, we have found that the main reason for the poor accuracy may lie in the extreme ill-conditioning of the matrices $A$ and $B$ in (10). From Table 2, where the condition numbers of the matrices $A$ and $B$ corresponding to different $M$ are shown, we can see that as the increase of $M$, the condition numbers of matrices $A$ and $B$ increase a lot. From Table 1 and Table 2, we believe that it is the increasingly ill-conditioned matrices $A$ and $B$ that prevent the solver from obtaining solutions with high accuracy.
Table 2: The condition number $\kappa(A)$ and $\kappa(B)$ obtained by Algorithm 2 with different $M$ for the 2D Laplace eigenvalue problem on a square domain $[0,1]^{2}$.
Dimension 100 200 300 400 500 600
$\kappa(A)$ 7.945e+15 2.383e+21 1.754e+26 8.161e+30 3.256e+33 6.145e+33
$\kappa(B)$ 2.954e+17 4.261e+23 2.024e+28 6.245e+32 8.485e+33 2.101e+34
4 Dimensionality reduced subspace method based on neural networks for eigenvalue problems
In the section above, we first simply apply the subspace method based on neural networks proposed in [49] to the solution of eigenvalue problems, and find that it behaves not so well. By some numerical tests, we find that the main reason is that as the dimension of the subspace $M$ increases, the condition number of the matrix $A$ and matrix $B$ become extremely large. In this case, problem (10) becomes extremely ill-conditioned, which results that solving (10) with a high precision becomes particularly difficult and almost impossible. A similar phenomenon has also been observed in random feature method [11], and the low-rank structure of random feature matrix have been systematically studied in [3, 7, 36, 47]. Therefore, in this section, we propose some strategy to reduce the condition number of the stiffness matrix and the mass matrix. In our opinion, the ill-conditioningness of the matrices $A$ and $B$ is resulted by the fact that the obtained basis functions are not completely linearly independent, leading to a lot of redundant information. Our basic idea is to find a way to remove the redundant functions, thereby reducing the dimension of the subspace but without significantly reducing the subspace information. Based on this idea, we apply some dimensionality reduction technique to the subspace spanned by the neural network based functions as introduced in Section 3, and then obtain an improved version of subspace method based on neural networks for eigenvalue problems, see Algorithm 3 for the details.
The only difference between Algorithm 2 and Algorithm 3 lies in that in Algorithm 3, a step to reduce the dimension of the subspace is added after obtaining the trained neural network basis functions. This step significantly reduces the condition number of the corresponding stiffness matrix and mass matrix, which can improve the solvability of the resulting algebraic eigenvalue problem and then improve the accuracy of the solution. After that, we calculate the Galerkin projection of the eigenvalue problem (5) onto this reduced subspace, and finally obtain an approximate solution.
Algorithm 3 General framework of the dimensionality reduced subspace method based on neural networks for eigenvalue problems
  1. Initialize the neural networks architecture (4), generate randomly $\theta$, and give $w_{1}$$\cdots$,$w_{k}$, then modify the neural network basis functions according to the boundary conditions and the computational domain.
  2. Update parameters $\theta$ by minimizing some loss function, and obtain the updated basis functions $\varphi_{1},\cdots,\varphi_{M}$.
  3. Reduce the dimension of the subspace $\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$ by some dimensionality reduction technique.
  4.Calculate the Galerkin projection of the eigenvalue problem (5) onto the reduced subspace and obtain an approximation of the solution.
4.1 Dimensionality reduction technique
In this subsection, we will introduce how we compress a large basis $\{\varphi_{1},\cdots,\varphi_{M}\}$ which numerically lacks good linear independece to a small basis $\{\psi_{1},\cdots,\psi_{K}\}$ with good linear independence while keeping as more necessary information as possible by some dimensionality reduction technique. Here, we choose the proper orthogonality decomposition (POD) to achieve this goal.
We can describe the problem as follows. Denote $V=\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$, $V_{k}\subset V$ any $K$-dimensional subspace of $V$, and $\Pi_{K}$ the orthogonal projection operator from $V$ to $V_{K}$. The problem is to find an optimal $K$-dimensional subspace $V_{K}^{\ast}$ to approximate the original space $V$ good enough. There will be two subproblems. One is that, for a given the dimension of the subspace $K\leq M$, how to find an optimal $K$-dimensional subspace $V_{K}^{\ast}$ to approximate original space $V$ on some least–squares sense, the other is how to choose suitable dimension of the subspace $K$. For the first problem, i.e. we seek an optimal orthogonal projection $\Pi_{K}^{\ast}:V\rightarrow V_{K}^{\ast}$ minimizing the projection error of $\bm{\Phi}=(\varphi_{1},\cdots,\varphi_{M})^{T}\in[L^{2}(\Omega)]^{M}$ under the following least–squares distance,
$\|\bm{\Phi}-\Pi_{K}\bm{\Phi}\|^{2}:=\sum\limits_{i=1}^{M}\big{|}\big{|}\varphi_{i}-\sum\limits_{j=1}^{K}(\varphi_{i},\psi_{j})\psi_{j}\big{|}\big{|}_{L^{2}(\Omega)}^{2}$
Hence, we need to solve the following minimization problem,
$\min_{\psi_{1},\cdots,\psi_{K}\in L^{2}(\Omega)}\sum\limits_{i=1}^{M}\big{|}\big{|}\varphi_{i}-\sum\limits_{j=1}^{K}(\varphi_{i},\psi_{j})\psi_{j}\big{|}\big{|}_{L^{2}(\Omega)}^{2}\ \text{s.t.}\ (\psi_{i},\psi_{j})=\delta_{ij},$ (13)
Denote $U=[u_{1},\cdots,u_{K}]\in\mathbb{R}^{M\times K}$ and $\bm{\Psi}=U^{T}\bm{\Phi}.$ We can rewrite the minimization problem (13) to the minimization problem (14) through following simple derivation,
$\sum\limits_{i=1}^{M}\big{|}\big{|}\varphi_{i}-\sum\limits_{j=1}^{K}(\varphi_{i},\psi_{j})\psi_{j}\big{|}\big{|}_{L^{2}(\Omega)}^{2}=\sum\limits_{i=1}^{M}\big{|}\big{|}\Phi^{T}e_{i}-\Phi^{T}\sum\limits_{j=1}^{K}\langle e_{i},u_{j}\rangle_{\mathcal{M}}u_{j}\big{|}\big{|}_{L^{2}(\Omega)}^{2}=\sum\limits_{i=1}^{M}\big{|}\big{|}e_{i}-\sum\limits_{j=1}^{K}\langle e_{i},u_{j}\rangle_{\mathcal{M}}u_{j}\big{|}\big{|}_{\mathcal{M}}^{2},$
where $\mathcal{M}=((\varphi_{i},\varphi_{j}))_{i,j}\in\mathbb{R}^{M\times M}$, $\langle x,y\rangle_{\mathcal{M}}=\langle x,\mathcal{M}y\rangle,\|x\|_{\mathcal{M}}=\sqrt{\langle x,x\rangle_{\mathcal{M}}},\forall x,y\in\mathbb{R}^{M}$ and $e_{1},\cdots,e_{M}$ are the standard basis in $\mathbb{R}^{M}$.
$\min_{u_{1},\cdots,u_{K}\in\mathbb{R}^{M}}\sum\limits_{i=1}^{M}\big{|}\big{|}e_{i}-\sum\limits_{j=1}^{K}\langle e_{i},u_{j}\rangle_{\mathcal{M}}u_{j}\big{|}\big{|}_{\mathcal{M}}^{2}\ \text{s.t.}\ \langle u_{i},u_{j}\rangle_{\mathcal{M}}=\delta_{ij},$ (14)
Notice that the problem (14) is nothing but the minimization problem considered in the POD method, which has been systematically studied (see [34, 42]), hence we give the following optimal orthogonal projection $\Pi_{K}^{\ast}$ of the problem (14) directly,
$\mathcal{M}^{2}u_{j}=\lambda_{j}\mathcal{M}u_{j},j=1,\cdots,K,$ (15)
$\min_{u_{1},\cdots,u_{K}\in\mathbb{R}^{M}}\sum\limits_{i=1}^{M}\big{|}\big{|}e_{i}-\sum\limits_{j=1}^{K}\langle e_{i},u_{j}\rangle_{\mathcal{M}}u_{j}\big{|}\big{|}_{\mathcal{M}}^{2}=\sum\limits_{j=K+1}^{M}\lambda_{j},$ (16)
where $\lambda_{1}\geq\cdots\lambda_{M}\geq 0$ are the first $M$ eigenvalues of $\mathcal{M}^{2}$ relative to $\mathcal{M}$, $u_{1},\cdots,u_{M}$ are the corresponding eigenvectors.
The vectors $u_{1},\cdots,u_{K}$ are called POD modes. In practical implementation, we can obtain them by solving the following symmetric $M\times M$ eigenvalue problem
$\mathcal{M}\bar{u}_{j}=\lambda_{j}\bar{u}_{j},j=1,\cdots,K,$
$u_{j}=\frac{1}{\sqrt{\lambda_{j}}}\bar{u}_{j},j=1,\cdots,K.$
For the second problem, since (16) shows the approximate error of the optimal $K$-dimensional subspace is related to the sum of the last $M-K$ eigenvalues of the matrix $\mathcal{M}$ (or the square root of the eigenvlaues), it’s reasonable to choose a small enough $K$ such that the first $K$ eigenvalues (or the square root of the eigenvlaues) account for a high proportion in the overall, i.e.,
$K=\mathop{\arg\min}\{I(N):I(N)\geq 1-\gamma\},$
where
$I(N)=\frac{\sum\limits_{j=1}^{N}\lambda_{j}}{\sum\limits_{j=1}^{M}\lambda_{j}}$ (17)
or
$I(N)=\frac{\sum\limits_{j=1}^{N}\sqrt{\lambda_{j}}}{\sum\limits_{j=1}^{M}\sqrt{\lambda_{j}}}$ (18)
and $\gamma$ is a given parameter.
In our experiment, we find that the indicator $I(N)$ defined by (18) works better.
We also have done a series of tests to see the change of condition numbers of stiffness matrix and mass matrix after dimensionality reduction operation. The condition numbers of stiffness matrix and mass matrix after the dimensionality reduction operation with $\gamma=10^{-9},10^{-10}$ and $10^{-11}$ are presented in Table 3. Comparing Table 3 with Table 2, we see clearly that the condition number of the matrices obtained before the dimensionality reduction operation increases rapidly with the increase of $M$. In contrast, the condition numbers of the matrices obtained after the dimensionality reduction operation can be controlled within an acceptable range and do not increase with the dimension of the subspace. Various numerical experiments have been done in Section 5. It can be observed that solving the eigenvalue problem in the reduced subspace indeed yields high-precision solutions of the eigenvalue problem (5) with an appropriate $\gamma$.
Table 3: The condition number $\kappa(\bar{A})$ and $\kappa(\bar{B})$ obtained by Algorithm 3 with different $M$ and dimensionality reduction parameter $\gamma$ for the 2D Laplace eigenvalue problem on a square domain $[0,1]^{2}$.
Condition number Dimension $\gamma=10^{-11}$ $\gamma=10^{-10}$ $\gamma=10^{-9}$
$\kappa(\bar{A})$ 100 1.00 1.00 1.00
200 1.00 1.00 1.00
300 1.00 1.00 1.00
400 1.00 1.00 1.00
500 1.00 1.00 1.00
600 1.00 1.00 1.00
$\kappa(\bar{B})$ 100 206.26 206.26 199.82
200 513.95 463.29 353.87
300 666.72 540.08 369.34
400 742.69 542.86 376.37
500 659.94 468.31 374.71
600 819.00 589.94 451.74
4.2 Galerkin projection of the original eigenvalue problem onto the reduced subspace
Now, we have obtained the reduced subspace $\text{span}\{\psi_{1},\psi_{2},\cdots,\psi_{K}\}$. We then calculate the Galerkin projection of the original eigenvalue problem (5) onto this subspace. From (3) in Section 2, the matrix eigenvalue problem can be formulated as:
$\bar{A}c=\lambda\bar{B}c,$ (19)
where $\bar{A}=(a(\psi_{i},\psi_{j}))_{ij},\bar{B}=(b(\psi_{i},\psi_{j}))_{ij}\in\mathbb{R}^{K\times K},c=(c^{(1)},\cdots,c^{(K)})^{T}\in\mathbb{R}^{K}.$
The first $k$ eigenpairs $(\lambda_{h,1},c_{1}),\cdots,(\lambda_{h,k},c_{k})$ can be obtained by solving the matrix eigenvalue problem (19) using some eigensolver. We then calculate the $w$ through $w_{i}=\tilde{R}c_{i},i=1,\cdots,k$, and finally obtain the approximation of the first $k$ eigenpairs $(k\leq K)$ $(\lambda_{h,1},u_{h,1}),\cdots,$ $(\lambda_{h,k},u_{h,k})$, where $u_{h,i}=\sum\limits_{j=1}^{M}w_{i}^{(j)}\varphi_{j}(x,\theta)$, $i=1,\cdots,k.$
Additionally, although the mass matrix $B$ in (10) is theoretically positive definite, when the $M$ is large, numerical errors during the generation of the mass matrix can lead to some very small negative eigenvalues. The dimensionality reduction operation effectively filters out these negative eigenvalues caused by numerical errors, making the mass matrix positive definite. Therefore, the resulted algebraic eigenvalue problem is easier to be solved with high accuracy.
5 Numerical experiments
In this section, we will use three typical eigenvalue problems, the Laplace eigenvalue problem, the decoupled harmonic oscillator and the coupled harmonic oscillator to demonstrate the performance of the methods proposed in our paper.
The algorithms are implemented based on the deep learning framework PyTorch. All the numerical experiments are carried on the LSSC-IV platform at the Academy of Mathematics and Systems Science, Chinese Academy of Sciences.
The following parameter settings are used in our numerical experiments: The neural networks architecture is implemented by a fully connected neural network with 3 hidden layers and a subspace layer with each hidden layer containing 100 neurons and the subspace layer containing $M$ neurons. We choose the activation function $\sigma(x)=\sin(x).$ During initialization part, the default initialization strategy is adopted to initialize neural network parameters $\theta$ and the random values generated according to a uniform distribution in the range [-1, 1] are applied to initialize coefficients $w_{1},\cdots,w_{k}$. The Adam optimizer with the default learning rate is used for training. In all experiments, we adopt the termination criteria introduced in (9) with the parameter $N_{max}=5000$. Different $\epsilon$ are chosen for different problems to achieve appropriate training (i.e., fewer epochs while maintaining accuracy). In the solving part, to improve the accuracy of the solution, we perform dimensionality reduction, with a given parameter denoted as $\gamma$. The solver $scipy.linalg.eigh$ in Python is used for solving the matrix eigenvalue problem (19). For reproducibility, all the random seeds used in the numerical experiment are set to 1.
We refer to the dimensionality reduced SNN for eigenvalue problems as R-SNN and compare our method with another neural-network based method proposed in [45], which can be viewed as a variant of DRM capable of solving multiple eigenpairs, simultaneously, denoted as DRM-M. For the fairness of the comparison, we choose the same neural networks architecture, initialization method, and the optimizer. R-SNN is trained until termination criterion (9) is satisfied, while DRM-M is trained for 25000 epochs.
The relative error of the approximate eigenvalues and eigenfunctions are defined as follows:
$err_{\lambda,l}:=\frac{|\lambda_{l}-\lambda_{h,l}|}{|\lambda_{l}|},err_{L^{2},l}:=\frac{\|u_{l}-u_{h,l}\|_{L^{2}(\Omega)}}{\|u_{l}\|_{L^{2}(\Omega)}},err_{H^{1},l}:=\frac{\|u_{l}-u_{h,l}\|_{H^{1}(\Omega)}}{\|u_{l}\|_{H^{1}(\Omega)}},$
where $\lambda_{l}$ and $u_{l}$ are reference eigenvalues and eigenfucntions, $l=1,\cdots,k$.
5.1 Laplace eignevalue problem
[…]
$\begin{cases}-\Delta u=\lambda u&\text{in }\Omega,\\ u=0&\text{on }\partial\Omega,\end{cases}$
where $\Omega=[0,1]^{2}$. The exact eigenvalues are as follows:
$\lambda_{n_{1},n_{2}}=\pi(n_{1}^{2}+n_{2}^{2}),n_{1},n_{2}=1,2,\cdots,$
and the corresponding eigenfunctions are
$u_{n_{1},n_{2}}(x_{1},x_{2})=\sqrt{2}\sin\left(\pi n_{1}x_{1}\right)\sin\left(\pi n_{2}x_{2}\right),n_{1},n_{2}=1,2,\cdots.$
We solve the first $15$ eigenpairs of the 2D Laplace eigenvalue problem using R-SNN. Since $\Omega$ is bounded, we choose the Legendre-Gauss integration scheme with 32 integration points in each dimension. For training, we set $\epsilon=10^{-3}$. And for the dimensionality reduction part, we set $\gamma=10^{-11}$.
Figure 1: The dimension of the reduced subspace and the relative errors of the first 15 eigenvalues resulted by different $M$ for the 2D Laplace eigenvalue problem obtained by R-SNN.
The dimension of the reduced subspace and the relative errors of the first 15 eigenvalues with different initial dimension $M$ obtained by R-SNN are presented in Figure 2. It can be observed that when the dimension of the original subspace $M$ is small, the dimension of the reduced subspace increases significantly with the increase of $M$. At this stage, the relative errors of the first 15 eigenvalues rapidly decay to below $10^{-12}$. When $M$ reaches 300, the increase in the dimension of the reduced subspace slows down significantly and the relative errors of the first 15 eigenvalues remain almost unchanged. Therefore, we only show results obtained by setting $M=300$ for both R-SNN and DRM-M.
Table 4: The relative errors of the first 15 eigenvalues with $M$=300 for the 2D Laplace eigenvalue problem.
Method $k$ ($n_{1},n_{2}$) $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$
R-SNN 1 (1,1) 3.960e-15 2.799e-08 3.912e-07
2 (2,1) 4.996e-14 4.711e-08 4.239e-07
3 (1,2) 1.526e-13 5.802e-08 5.366e-07
4 (2,2) 1.449e-13 6.954e-08 5.151e-07
5 (3,1) 1.523e-13 8.539e-08 5.753e-07
6 (1,3) 1.892e-13 9.419e-08 6.128e-07
7 (3,2) 1.256e-13 1.137e-07 6.993e-07
8 (2,3) 3.792e-13 1.284e-07 7.331e-07
9 (4,1) 9.622e-14 1.416e-07 7.757e-07
10 (1,4) 2.531e-13 1.436e-07 7.976e-07
11 (3,3) 1.923e-13 1.774e-07 9.253e-07
12 (4,2) 1.518e-13 1.737e-07 8.788e-07
13 (2,4) 5.876e-13 2.258e-07 1.181e-06
14 (4,3) 2.919e-13 2.558e-07 1.096e-06
15 (3,4) 4.750e-13 3.231e-07 1.435e-06
DRM-M 1 (1,1) 3.868e-07 1.222e-04 7.086e-04
2 (2,1) 1.325e-07 1.135e-04 4.374e-04
3 (1,2) 2.186e-07 1.249e-04 6.110e-04
4 (2,2) 7.583e-08 9.439e-05 3.828e-04
5 (3,1) 1.126e-07 2.027e-04 5.936e-04
6 (1,3) 1.608e-07 1.709e-04 4.993e-04
7 (3,2) 8.341e-08 1.899e-04 4.586e-04
8 (2,3) 1.163e-07 1.825e-04 4.520e-04
9 (4,1) 3.046e-07 1.360e-03 1.962e-03
10 (1,4) 1.193e-06 5.145e-04 9.427e-04
11 (3,3) 1.106e-07 2.519e-04 4.826e-04
12 (4,2) 2.862e-07 1.372e-03 1.908e-03
13 (2,4) 1.039e-06 5.832e-04 9.282e-04
14 (4,3) 3.565e-07 1.184e-03 1.516e-03
15 (3,4) 6.469e-07 8.103e-04 1.084e-03
Table 5: The epochs resulted by different $M$ for the 2D Laplace eigenvalue problem obtained by R-SNN.
Dimension 50 100 150 200 250 300 350 400 450 500 550 600
Epoch 177 195 190 342 160 165 171 160 271 168 200 191
The relative errors of the first 15 eigenpairs obtained using R-SNN and DRM-M with the dimension of original subspace $M=300$ are presented in Table 4, and epochs required for our training process are shown in Table 5. It can be seen that the accuracy of DRM-M reaches only $10^{-8}$ for eigenvalues and $10^{-4}$ for eigenfunctions after 25000 epochs, while our method achieves accuracy below $10^{-12}$ for eigenvalues and nearly $10^{-7}$ for eigenfunctions, only with 100-400 epochs.
5.2 Harmonic oscillator problems
In this subsection, we will study the following 2D decoupled and coupled harmonic oscillator.
5.2.1 Two-dimensional decoupled harmonic oscillator
[…]
$-\frac{1}{2}\Delta u+Vu=\lambda u\quad\text{in}\quad\Omega,$
where $\Omega=\mathbb{R}^{2},V(x_{1},x_{2})=\frac{1}{2}(x_{1}^{2}+x_{2}^{2})$. The exact eigenvalues are given as follows:
$\lambda_{n_{1},n_{2}}=(\frac{1}{2}+n_{1})+(\frac{1}{2}+n_{2}),n_{1},n_{2}=0,1,\cdots,$
and the corresponding eigenfunctions are
$u_{n_{1},n_{2}}(x_{1},x_{2})=\mathcal{H}_{n_{1}}(x_{1})e^{-x_{1}^{2}/2}\mathcal{H}_{n_{2}}(x_{2})e^{-x_{2}^{2}/2},n_{1},n_{2}=0,1,\cdots,$
where $\mathcal{H}_{n}$ is Hermite polynomials [4]. We solve the first $15$ eigenpairs of 2D decoupled harmonic oscillator using R-SNN. Since $\Omega$ is $\mathbb{R}^{2}$, the Hermite-Gauss integration scheme is chosen to compute the unbounded integral, with 99 integration points in each dimension. In training part, we take $\epsilon=10^{-2}$. In the dimensionality reduction part, we take $\gamma=10^{-10}$.
The dimension of the reduced subspace and the relative errors of the first 15 eigenvalues under different initial dimension $M$ are presented in Figure 3. It can be observed that when the original dimension of the subspace is small, the dimension of the reduced subspace increases significantly with the increase of $M$. At this stage, the relative errors of the first 15 eigenvalues rapidly decay to below $10^{-11}$. When $M$ reaches 900, the increase in the dimension of the reduced subspace slows down significantly and the relative errors of the first 15 eigenvalues remain almost unchanged. Therefore, we only show results obtained by setting $M=900$ for both R-SNN and DRM-M.
Figure 2: The dimension of the reduced subspace and the relative errors resulted by different $M$ for 2D decoupled harmonic oscillator obtained by R-SNN.
Table 6: The relative errors of the first 15 eigenvalues with $M$=900 for 2D decoupled harmonic oscillator.
Method $k$ $(n_{1},n_{2})$ $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$
R-SNN 1 (0,0) 7.294e-14 3.977e-08 2.373e-07
2 (0,1) 8.005e-14 6.304e-08 3.012e-07
3 (1,0) 8.926e-14 6.095e-08 2.963e-07
4 (0,2) 3.449e-14 1.184e-07 4.473e-07
5 (1,1) 4.737e-14 9.585e-08 3.648e-07
6 (2,0) 1.198e-13 9.863e-08 4.040e-07
7 (0,3) 7.283e-14 3.044e-07 1.035e-06
8 (1,2) 9.104e-14 1.458e-07 5.599e-07
9 (2,1) 2.474e-13 1.632e-07 5.801e-07
10 (3,0) 7.412e-13 2.373e-07 8.030e-07
11 (1,3) 3.457e-12 1.312e-06 3.803e-06
12 (2,2) 4.765e-12 9.109e-07 2.550e-06
13 (3,1) 5.116e-12 1.221e-06 3.586e-06
14 (0,4) 1.962e-11 1.923e-06 5.044e-06
15 (4,0) 2.506e-11 1.185e-06 3.308e-06
DRM-M 1 (0,0) 2.449e-09 2.062e-05 3.913e-05
2 (0,1) 4.052e-09 3.637e-05 6.912e-05
3 (1,0) 8.495e-09 5.009e-05 9.015e-05
4 (0,2) 3.324e-09 4.887e-05 7.840e-05
5 (1,1) 4.008e-09 5.239e-05 8.106e-05
6 (2,0) 6.147e-09 4.121e-05 6.859e-05
7 (0,3) 3.407e-09 7.011e-05 1.035e-04
8 (1,2) 3.609e-09 6.003e-05 8.251e-05
9 (2,1) 7.057e-09 6.483e-05 9.069e-05
10 (3,0) 8.040e-09 7.635e-05 1.054e-04
11 (1,3) 6.594e-09 1.537e-04 1.734e-04
12 (2,2) 7.201e-09 1.784e-04 2.085e-04
13 (3,1) 1.029e-08 1.329e-04 1.632e-04
14 (0,4) 1.426e-08 1.176e-04 1.424e-04
15 (4,0) 2.001e-08 1.511e-04 1.772e-04
Table 7: The epochs resulted by different $M$ for 2D decoupled harmonic oscillator obtained by R-SNN.
Dimension 100 200 300 400 500 600 700 800 900 1200 1500 1800
Epoch 20 21 23 23 22 20 20 21 21 21 20 22
The relative errors of the first 15 eigenpairs obtained using R-SNN and DRM-M with the dimension of the original subspace $M=900$ are presented in Table 6. Table 7 provides the number of epochs for various $M$. Here, it is evident that our accuracy is significantly superior to the results of DRM-M with far fewer epochs. Additionally, results from other studies for the same example are also considered. An accuracy of $10^{-2}$ is achieved using machine learning methods based on Monte Carlo integration, as demonstrated in [28]. Furthermore, our accuracy surpasses that of [45], which employs numerical integration based on TNN to achieve high-precision calculations. Compared to [45], which requires 500,000 epochs, our method achieves the desired accuracy with only 20 epochs.
5.2.2 Two-dimensional coupled harmonic oscillator
[…]
$-\frac{1}{2}\Delta u+Vu=\lambda u\quad\text{in}\quad\Omega,$
where $\Omega=\mathbb{R}^{2},V(x_{1},x_{2})=\frac{1}{2}(a_{11}x_{1}^{2}+2a_{12}x_{1}x_{2}+x_{22}^{2})$, the coefficients $a_{11},a_{12},a_{22}$ are selected from [28] and maintains the same accuracy as [28], i.e. $a_{11}=0.8851,a_{12}=-0.1382,a_{22}=1.1933$.
The coupled harmonic oscillator defined by the above potential energy term can be decoupled into the following diagonal form:
$V(y_{1},y_{2})=\frac{1}{2}(\mu_{1}y_{1}^{2}+\mu_{2}y_{2}^{2}),$
where $y_{1}=-0.9339352418x_{1}-0.3574422527x_{2}$ and $y_{2}=0.3574422527x_{1}-0.93393524$ $18x_{2}$, $\mu_{1}=0.8322071257$, $\mu_{2}=1.2461928742$.
Based on the decoupled form, the exact eigenvalues are given as follows:
$\lambda_{n_{1},n_{2}}=(\frac{1}{2}+n_{1})\mu_{1}^{1/2}+(\frac{1}{2}+n_{2})\mu_{2}^{1/2},n_{1},n_{2}=0,1,\cdots,$
and the corresponding eigenfunctions are
$u_{n_{1},n_{2}}(x_{1},x_{2})=\mathcal{H}_{n_{1}}(\mu_{1}^{1/4}y_{1})e^{-\mu_{1}^{1/2}y_{1}^{2}/2}\mathcal{H}_{n_{2}}(\mu_{2}^{1/4}y_{2})e^{-\mu_{2}^{1/2}y_{2}^{2}/2},n_{1},n_{2}=0,1,\cdots.$
We solve the first $15$ eigenpairs of 2D coupled harmonic oscillator using R-SNN. Since $\Omega$ is $\mathbb{R}^{2}$, we choose the Hermite-Gauss integration scheme, with 99 integration points in each dimension. In training part, we take $\epsilon=10^{-2}$. In the dimensionality reduction part, we take $\gamma=10^{-10}$.
Figure 3: The dimension of the reduced subspace and the relative errors resulted by different $M$ for 2D coupled harmonic oscillator obtained by R-SNN.
The dimension of the reduced subspace and the relative errors of the first 15 eigenvalues under different initial dimensions $M$ are presented in Figure 4. A numerical performance similar to decoupled harmonic oscillator can be observed. Accordingly, we present specific results for $M=900$ for both R-SNN and DRM-M.
Table 8: The relative errors of the first 15 eigenvalues with $M$=900 for 2D coupled harmonic oscillator.
Method $k$ $(n_{1},n_{2})$ $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$
R-SNN 1 (0,0) 2.330e-11 4.430e-08 2.520e-07
2 (1,0) 1.979e-11 6.415e-08 3.183e-07
3 (0,1) 2.647e-11 7.323e-08 3.517e-07
4 (2,0) 1.939e-11 2.710e-07 1.011e-06
5 (1,1) 2.348e-11 1.704e-07 6.191e-07
6 (0,2) 2.740e-11 1.188e-07 4.584e-07
7 (3,0) 2.829e-11 1.107e-06 3.355e-06
8 (2,1) 2.312e-11 4.443e-07 1.350e-06
9 (1,2) 2.545e-11 2.514e-07 8.112e-07
10 (0,3) 2.798e-11 1.478e-07 5.192e-07
11 (4,0) 2.524e-10 6.010e-06 1.645e-05
12 (3,1) 1.159e-10 4.032e-06 1.036e-05
13 (2,2) 4.021e-11 1.626e-06 4.390e-06
14 (1,3) 3.324e-11 1.120e-06 2.951e-06
15 (0,4) 2.874e-11 3.594e-07 1.063e-06
DRM-M 1 (0,0) 1.154e-07 1.361e-04 2.723e-04
2 (1,0) 2.964e-08 9.442e-05 1.736e-04
3 (0,1) 2.188e-08 9.079e-05 1.600e-04
4 (2,0) 6.870e-08 1.912e-04 2.966e-04
5 (1,1) 1.830e-08 9.054e-05 1.414e-04
6 (0,2) 1.826e-08 1.172e-04 1.754e-04
7 (3,0) 5.469e-08 1.966e-04 2.674e-04
8 (2,1) 4.192e-08 2.110e-04 2.704e-04
9 (1,2) 1.655e-08 1.173e-04 1.616e-04
10 (0,3) 1.759e-08 1.341e-04 1.743e-04
11 (4,0) 9.014e-08 3.495e-04 4.336e-04
12 (3,1) 6.991e-08 3.703e-04 4.364e-04
13 (2,2) 3.048e-08 2.252e-04 2.719e-04
14 (1,3) 4.191e-08 2.746e-04 3.394e-04
15 (0,4) 3.972e-08 2.894e-04 3.600e-04
Table 9: The epochs resulted by different $M$ for 2D coupled harmonic oscillator obtained by R-SNN.
Dimension 100 200 300 400 500 600 700 800 900 1200 1500 1800
Epoch 20 21 22 23 22 20 20 21 20 21 20 23
The relative errors of the first 15 eigenpairs obtained from R-SNN and DRM-M for solving the 2D coupled harmonic oscillator with dimension of the original subspace $M=900$ are shown in Table 8. Similar to the decoupled harmonic oscillator, our method provides solutions with higher accuracy while requiring fewer epochs during training, as shown in Table 9, compared to the DRM-M and other existing results in [28] and [45].
6 Conclusions
In this paper, we propose a subspace method based on neural networks for eigenvalue problems with high accuracy. SNN can in fact be viewed as a method combining both the machine learning method with the traditional methods, that is, use the machine learning method to get some finite dimensional subspace, then calculate the Galerkin projection of the original eigenvalue problems onto the subspace and obtain a approximate solution, just like what the traditional methods do. Through analysis and experimental validation, we find that reducing the condition number of the stiffness matrix and the mass matrix through a dimensionality reduction technique can effectively improve the accuracy of approximate solution. Our method achieves high precision with only a few training epochs. The eigenvalue errors of our method can be below $10^{-11}$, and the eigenfunction errors are nearly below $10^{-7}$, with the required number of epochs ranging from 20 to 400.
Acknowledgments
We would like to acknowledge Professor Aihui Zhou for his helpful discussions that improve the presentation of this paper.

### v2

@@
-
+Subspace method based on neural networks for solving eigenvalue problems  [^1] This work was supported by the National Natural Science Foundation of China under grants 92270206 and 12071045 and the Fund of National Key Laboratory of Computational Physics
 Abstract
-With the rapid development of machine learning, numerical discretization methods based on deep neural networks have been widely used in many fields, especially in solving high-dimensional problems where traditional methods face bottlenecks.
-However, for low-dimensional problems, existing machine learning methods are not as accurate and efficient as traditional methods.
 In this paper, we propose a subspace method based on neural networks for eigenvalue problems with high accuracy and low cost.
-Our basic idea is to use neural network based basis functions to span a subspace, then calculate the parameters of the neural network based basis functions through appropriate training, and finally calculate the Galerkin projection of the eigenvalue problem onto the subspace and obtain an approximate solution.
-In addition, we reduce the dimension of the subspace by applying some dimensionality reduction technique, which can improve the accuracy of the approximate solution in further.
-Numerical experiments show that we can obtain approximate eigenvalues with accuracy of $10^{-11}$ but with less then 400 epochs, which is significantly superior to other existing neural network based methods.
+We first construct a neural network-based orthogonal basis by some deep learning method and dimensionality reduction technique, and then calculate the Galerkin projection of the eigenvalue problem onto the subspace spanned by the orthogonal basis and obtain an approximate solution.
+Numerical experiments show that we can obtain approximate eigenvalues and eigenfunctions with very high accuracy but low cost.
 keywords:
-subspace, basis function, eigenvalue problem, multi-eigenpairs, neural networks, high-accuracy
+eigenvalue problems, subspace, method neural networks orthogonal basis, Galerkin projection
 1 Introduction
@@
 Over recent decades, many numerical methods have been developed to discretize eigenvalue problems with great success, including the finite difference method [29], the finite volume method [16, 35], the finite element method [6, 10, 48], and the spectral method [22].
-With the rapid development of machine learning, many numerical discretization methods for partial differential equations (PDEs) based on deep neural networks have received increasing attention, such as the physical-information neural networks (PINN) [35], the deep Galerkin method (DGM) [39], the deep Ritz method (DRM) [50], the deep Nitsche method (DNM) [32], and the weak adversarial networks (WAN) [52].
-By these methods, the solution of the partial differential equation is represented by a neural network directly and the optimal parameters of the neural network is obtained by minimizing a loss function.
+With the rapid development of machine learning, many numerical discretization methods for partial differential equations (PDEs) based on deep neural networks have received increasing attention, such as the physical-information neural networks (PINN) [40], the deep Galerkin method (DGM) [47], the deep Ritz method (DRM) [58], and the weak adversarial networks (WAN) [60].
+By these methods, the solution of a partial differential equation is represented by neural networks directly and the optimal parameters of neural networks are obtained by minimizing some loss function.
 The main difference between these methods lies in the choice of the loss function.
 For instance, PINN and DGM use the $L^{2}$ norm of the residual in the strong form as the loss function, while DRM and WAN use the energy functional and the operator norm in the weak form as the loss function, respectively.
-In contrast, DNM uses a loss function based on Nitsche’s energy formulation to handle essential boundary conditions.
-Due to the powerful approximation capabilities of neural networks, the aforementioned methods are not only used to deal with source problems but also to handle eigenvalue problems.
+Due to the powerful approximation capabilities of neural networks [4, 18, 25, 26, 30, 37, 38, 44, 45, 46], the aforementioned methods are used to deal with many different kinds of partial differential equations, including not only some source problems but also some eigenvalue problems.
 For instance, in [24], a wave function ansatz based on deep neural networks is introduced to solve the many-electron Schrödinger equation.
-In [23, 24], various wave function ansatzes in the form of a Slater determinant are also introduced for the Schrödinger equation.
-In [28], a neural network based multistate solver is proposed for finding the first $k$ eigenpairs of the Schrödinger equation simultaneously.
+In [Hermann_Schätzle_Noé_2020, Hermann_Spencer_Choo_Mezzacapo_Foulkes_Pfau_Carleo_Noé_2023], various wave function ansatzes in the form of a Slater determinant are also introduced for the Schrödinger equation.
+In [32], a neural network-based multistate solver is proposed for finding the first $k$ eigenpairs of the Schrödinger equation simultaneously.
 In [23], a diffusion Monte Carlo-like method is proposed for handling high-dimensional eigenvalue problems.
 In [31], a semigroup method is provided for solving high-dimensional elliptic partial differential equations using neural networks and is extended to a primal-dual algorithm for solving eigenvalue problems.
-In [29], a type of tensor neural network based on the DECOMP/PARAFAC (CP) tensor decomposition is proposed, along with an efficient numerical integration scheme for tensor-type functions.
+In [33], a type of tensor neural network-based on the DECOMP/PARAFAC (CP) tensor decomposition is proposed, along with an efficient numerical integration scheme for tensor-type functions.
 Theoretical analyses and applications of tensor neural networks are provided in [52] and [51, 53], respectively.
-For example, tensor neural networks have been used to solve the the first eigenvalue and the corresponding eigenfunction of Schrödinger equation in [43] and the first $k$ eigenpairs of high-dimensional eigenvalue problems in [45].
-Additionally, in [51], a machine learning method is proposed for computing eigenvalues and eigenfunctions of the Schrödinger equation, which employs trial functions satisfying the homogeneous boundary conditions to avoid the penalty of the boundary, and an explicit convergence rate of the generalization error also is provided in the article.
-While deep neural network based methods have made strides in solving high-dimensional problems where traditional methods fail, they face challenges in solving low-dimensional problems due to the lack of algorithms to train neural networks efficiently.
-As a result, they lag behind traditional methods in accuracy and efficiency, which prevent most deep neural network based methods from being not only usable but also efficient.
+For example, tensor neural networks have been used to solve the smallest eigenvalue and the corresponding eigenfunction of Schrödinger equation in [51] and the first $k$ eigenpairs of high-dimensional eigenvalue problems in [53].
+Additionally, in [59], a machine learning method is proposed for computing eigenvalues and eigenfunctions of the Schrödinger equation, and an explicit convergence rate of the generalization error is also provided in the article.
+While deep neural network-based methods have made strides in solving high dimensional problems where classical methods fail, they face challenges in solving low-dimensional problems due to the lack of algorithms to train neural networks efficiently.
+As a result, they lag behind classical methods in accuracy and efficiency, which prevent most deep neural network-based methods from being not only usable but also efficient.
 In recent years, some emerging methods based on shallow neural networks have received increasing attention, such as the extreme learning machine (ELM) [27] and the random feature method (RFM) [11, 14].
-In these methods, all parameters of the neural network except for the output layer are randomly generated and fixed, which simplifies the optimization problem into a linear system.
+In these methods, all parameters of neural networks except for the output layer are randomly generated and fixed, which simplifies the optimization problem into a linear system.
 By using the least-squares method to solve the linear system, the expensive training process can be carried out at a low cost with even higher accuracy.
 These methods and their variants [20, 43, 49] have been widely used for solving partial differential equations [11, 13, 34, 54].
-However, as so far, there is little application of these methods for solving eigenvalue problems.
+However, so far, there is little application of these methods for solving eigenvalue problems.
 From another perspective, the methods above fundamentally offer a method to construct basis functions using neural networks.
 However, as the basis functions constructed by neural networks are sensitive to the choice of hyperparameters, improper choices can significantly impact the accuracy of the approximate solution.
-Randomly generating these hyperparameters is proved to be inefficient for some problems, which means that the search for more reasonable ways to select and update hyperparameters is a worthwhile pursuit.
+Randomly generating these hyperparameters may be inefficient for some problems, which means that the search for more reasonable ways to select and update hyperparameters is a worthwhile pursuit.
 To address the impact of hyperparameter selection on solutions, a modified batch intrinsic plasticity method for pre-training the randomly generated hyperparameters in ELM is proposed in [19].
-A differential evolution algorithm is employed to compute the optimal or near-optimal parameter $R_{m}$, where the parameters of the neural network are set to random values generated in interval $[-R_{m},R_{m}]$ in [19].
+A differential evolution algorithm is employed to compute the optimal or near-optimal parameter $R_{m}$, where the parameters of the neural networks are initialized with random values generated in the interval $R_{m},R_{m}]$ in [21].
 Additionally, a better way to randomly select hyperparameters based on some geometric explanations is provided in [61].
-There are also some methods that couple the updating of hyperparameters with the solution of the linear system.
-An alternative iteration between gradient descent to update the hyperparameters and least squares to solve the linear system is adopted in [2, 14].
-Recently, a subspace method based on neural networks (SNN) for partial differential equations in strong form has been proposed in [49], which provides a deterministic way to obtain the appropriate hyperparameters.
-The main idea of the method is to use neural network based functions as basis functions to span a subspace, then obtain the parameters of the neural network based functions by minimizing a loss function and subsequently find the approximate solution in the subspace.
-Unlike ELM-like methods and random feature methods, parameters of the neural network in SNN are not randomly generated; instead, they are obtained by minimizing some loss function.
-In comparison to classical deep neural network based discretization methods, the training process in the SNN focuses on finding better basis functions, which requires relatively lower accuracy in the training process and often needs significantly fewer training epochs.
+There are also some deterministic algorithms to calculate hyperparameters.
+An alternating iteration process is introduced in [2, 15], which combines gradient descent for updating the hyperparameters and least squares for solving the linear system.
+A subspace method based on neural networks (SNN) for partial differential equations has been proposed in [57].
+The main idea of this method is to train neural network-based functions by minimizing a loss function, and then use least squares to find the approximation of the original problem in the subspace spanned by those trained functions, with complete decoupling between training the subspace and solving the linear system.
+Unlike ELM-like methods and random feature methods, the parameters of the neural network-based functions in SNN are not randomly generated; instead, they are obtained by minimizing some loss function.
+In comparison to classical deep neural network-based discretization methods, the training process in the SNN focuses on finding better basis functions, which requires relatively lower accuracy in the training process and often needs significantly fewer training epochs.
+Compared with methods updating hyperparameters by alternating iteration, the updating of hyperparameters and solution of the linear system in SNN is totally decoupled, which greatly reduces the computational cost.
 Consequently, SNN can achieve high accuracy with low cost in training for solving partial differential equations.
 Furthermore, a weak form version of SNN has also been proposed in [36].
-From another perspective, the training process of SNN essentially offers a deterministic approach to generate a basis based on neural networks.
-The solving process of SNN, which calculate the Galerkin projection of the original problem onto the subspace spanned by the basis functions obtained from the training process, can be viewed as a special case of finite dimensional approximation commonly employed in traditional methods.
-The difference lies in the selection of basis functions: traditional methods like the finite element method employs piecewise polynomials as basis functions, spectral method employs orthogonal polynomials as basis functions, whereas SNN employs appropriately trained neural network based functions as basis functions.
-In this paper, motivated by the work in [49], we first propose a subspace method based on neural networks for eigenvalue problems and then an improved version.
-In the improved version, we add a dimensionality reduction step after obtaining the subspace spanned by the neural network based basis functions.
-Our method mainly consist of four steps.
-First, we specify neural networks architecture and appropriately modify basis functions as needed for the specific problem.
-Next, we obtain the parameters of the basis functions by minimizing some loss function.
-Then, we reduce the dimension of the subspace spanned by the basis functions by some dimensionality reduction technique.
-Finally, we calculate the Galerkin projection of the original eigenvalue problem onto the reduced subspace and obtain an approximate solution.
-We apply our method to solution of some typical eigenvalue problems.
-In those examples, our algorithm can achieve very high accuracy with few training epochs, which demonstrates the advantages of our new algorithm in both computational cost and accuracy.
+In another view, the training process of SNN essentially offers a deterministic approach to update the parameters of the neural network-based functions.
+The solving process of SNN, which finds an approximate solution in the subspace spanned by the trained functions, can be viewed as a kind of finite dimensional approximation commonly employed in classical methods.
+The difference lies in the selection of basis functions: classical methods like the finite element method employ piecewise polynomials as basis functions, spectral method employs orthogonal polynomials as basis functions, whereas SNN employs trained neural network-based functions as basis functions.
+In this paper, motivated by the idea of using deep learning method to train the basis functions introduced in [57], we propose a subspace method based on neural networks for eigenvalue problems.
+Besides using the deep learning method to train some neural network-based functions, we further introduce some dimensionality reduction technique to the trained functions and obtain some neural network-based orthogonal basis.
+Our method mainly consists of two steps.
+First, we construct a neural network-based orthogonal basis by some deep learning method and dimensionality reduction technique.
+Then, we use some classical method, e.g., the Galerkin method, to discretize the eigenvalue problem in the subspace spanned by the orthogonal basis, and then obtain the approximate solution.
+We apply our method to the solution of some typical eigenvalue problems.
+The numerical results show that our algorithm can achieve very high accuracy with few training epochs, which demonstrates the advantages of our new algorithm in both computational cost and accuracy.
 The subsequent parts of our paper are organized as follows.
-In Section 2, we provide some preliminary knowledge, including some notation and classical results for a linear elliptic eigenvalue problem, and an briefly introduction of the subspace method based on neural networks for some source problems.
-In Sections 3, we introduce the general framework of the subspace method based on the neural networks for eigenvalue problems and the details for each step.
-In Section 4, we introduce some dimensionality reduction technique to reduce the dimension of the subspace spanned by the neural network based functions, and obtain the improved version of SNN for eigenvalue problems.
+In Section 2, we provide some preliminary knowledge, including some notation and classical results for a linear elliptic eigenvalue problem, and a brief introduction of the subspace method based on neural networks for some source problems.
+In Sections 3, we introduce the general framework of our subspace method based on the neural networks for eigenvalue problems and the details for each step,especially on how to construct neural network-based orthogonal basis.
 In Section 4, we provide some numerical results to show the efficiency of our new method.
@@
 2 Preliminaries
-Before introducing our method, we first give some background knowledge, including some basic results of a elliptic eigenvalue problem and some briefly introduction to the subspace method based on neural networks for partial differential equations proposed in [49].
+Before introducing our method, we first give some preliminary knowledge, including some basic results of a elliptic eigenvalue problem and a brief introduction to the subspace method based on neural networks for partial differential equations proposed in [57].
 2.1 A linear elliptic eigenvalue problem and its finite dimensional discretization
-[…]
+Let ${\rm\Omega}\subset\mathbb{R}^{d}(d\geq 1)$ be a polygonal domain.
+We shall use the standard notation for Sobolev spaces $W^{s,p}(\rm\Omega)$ and their associated norms (see, e.g.
+[1]).
+We denote $H_{0}^{1}({\rm\Omega})=\{v\in H^{1}({\rm\Omega}):v|_{\partial{\rm\Omega}}=0\}$, $(\cdot,\cdot)$ be the standard $L^{2}$ inner product and $\langle\cdot,\cdot\rangle$ be the standard Euclidean inner product.
+Additionally, in this paper, we will use the letter $C$ to denote a generic positive constant which may stand for different values at its different occurrences.
+For convenience, following [56], the symbols $\lesssim$ are used in this paper, where $x\lesssim y$ means that $x\leq Cy$ for some constant $C$ that is independent of the spatial discretization.
 Consider the following elliptic eigenvalue problem: Find $(\lambda,u)\in\mathbb{R}\times V$ such that $b(u,u)=1$ and
 $a(u,v)=\lambda b(u,v),\quad\forall v\in V.$ (1)
-[…]
+Here, $(V,\|\cdot\|)$ is a Hilbert space with inner product ($\cdot$$\cdot$) and norm $\|\cdot\|$. The bilinear forms $a(\cdot,\cdot)$ and $b(\cdot,\cdot)$ on $V\times V$ are symmetric and satisfy
 $a(u,v)\lesssim\|u\|\ \|v\|,\forall u,v\in V,$
@@
 $\|\cdot\|_{a}=\sqrt{a(\cdot,\cdot)},\|\cdot\|_{b}=\sqrt{b(\cdot,\cdot)}.$
-[…]
+Suppose the norm $\|\cdot\|_{b}$ is compact with respect to $\|\cdot\|$ ([7]), i.e., from any sequence which is bounded in $\|\cdot\|$, one can extract a subsequence which is Cauchy in $\|\cdot\|_{b}$.
 It is well known that the eigenvalue problem (1) has an eigenvalue sequence $\{\lambda_{k}\}$ ([6])
@@
 $a(u_{h},v_{h})=\lambda b(u_{h},v_{h}),\quad\forall v_{h}\in V_{h}.$ (2)
-Suppose $V_{h}=\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$, where $\varphi_{1},\cdots,\varphi_{M}$ are basis functions of the finite dimensional subspace $V_{h}$.
+Suppose $V_{h}=\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$, where $\{\varphi_{i}\}_{i=1}^{M}$ is a basis of the finite dimensional subspace $V_{h}$.
 Substituting $u_{h}=\sum_{j=1}^{M}w^{(j)}\varphi_{j}\ \text{and choosing}\ v_{h}\ \text{to be}\ \varphi_{i}$ into (2), we then have:
@@
 Denote $A$ the stiffness matrix and $B$ the mass matrix, that is, $A=(a(\varphi_{i},\varphi_{j}))_{ij}\in\mathbb{R}^{M\times M},B=(b(\varphi_{i},\varphi_{j}))_{ij}\in\mathbb{R}^{M\times M}$.
-The eigenvalue problem (1) can be formulated as the following matrix eigenvalue problem:
-$Aw=\lambda Bw,$ (3)
-where $w=(w^{(1)},\cdots,w^{(M)})^{T}\in\mathbb{R}^{M}.$
-By solving the matrix eigenvalue problem (3), an approximate solution to the eigenvalue problem (1) can be obtained.
+The eigenvalue problem (1) can be formulated as the following algebraic eigenvalue problem:
+$A\bm{w}=\lambda B\bm{w},$ (3)
+where $\bm{w}=(w^{(1)},\cdots,w^{(M)})^{T}\in\mathbb{R}^{M}.$
+By solving the algebraic eigenvalue problem (3), an approximate solution to the eigenvalue problem (1) is then obtained.
 2.2 Subspace method based on neural networks
 In this subsection, we will give a brief review for the subspace method based on neural networks proposed in [57] for the partial differential equations, which has been proven to be able to achieve high accuracy with low cost.
+In [57], the following general source problem is considered:
+$\begin{cases}\mathscr{A}u=f&\text{in }\rm\Omega,\\ \mathscr{B}u=g&\text{on }\partial\rm\Omega,\end{cases}$
+where $\mathscr{A}$ and $\mathscr{B}$ are two differential operators, $f$ and $g$ are given functions.
 The neural networks architecture used in [57] is comprised of four parts: an input layer, several hidden layers, a subspace layer and an output layer.
-For simplicity, we take the case of single output as an example to show the mathematical expression of the neural network.
+For simplicity, we take the case of single output as an example to show the mathematical expression of the neural networks.
 The expression of multiple outputs can be obtained similarly.
-Denote $L$ the number of hidden layers, $n_{0}=d$ the number of neurons in the input layer, $n_{1},\cdots,n_{L}$ the number of neurons in each hidden layer, $n_{L+1}=M$ the number of neurons in the subspace layer, and $n_{L+2}=1$ the number of neurons in the output layer.
-The neural network used in [49] can be expressed as follows:
+Denote $L$ the number of hidden layers, $n_{0}=d$ the number of neurons in the input layer, $n_{1},\cdots,n_{L}$ the number of neurons in each hidden layer, $n_{L+1}=M$ the number of neurons in the subspace layer.
+The neural networks used in [57] can be expressed as follows:
 […]
-Here, $W_{l}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ and $\bm{b}_{l}\in\mathbb{R}^{n_{l}}$ are the weights and biases of the corresponding layer, respectively, $\bm{x}\in\mathbb{R}^{d}$ is the input, $\varphi=(\varphi_{1},\cdots,\varphi_{M})^{T}$ and $w=(w^{(1)},\cdots,w^{(M)})^{T}$, $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ is an activate function which can be extended to $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ as follows:
-$\sigma(\bm{x})=\begin{bmatrix}\sigma(x_{1})\\ \vdots\\ \sigma(x_{n})\end{bmatrix}\ \text{for any}\ \bm{x}=(x_{1},\cdots,x_{n})^{T}\ \in\mathbb{R}^{n}.$
-Denote $\theta=\{W_{1},\cdots,W_{L+1},\bm{b}_{1},\cdots,\bm{b}_{L+1}\}\in\Theta:=\mathbb{R}^{n_{1}\times n_{0}}\times\cdots\times\mathbb{R}^{n_{L+1}\times n_{L}}\times\mathbb{R}^{n_{1}}\times\cdots\times\mathbb{R}^{n_{L+1}}$ the parameters in the neural network, then $u(\bm{x};\theta,w)$ is the output of $\bm{x}$ with respect to parameters $\theta$ and $w$.
+Here, $\mathcal{T}_{l}(\bm{y})=W_{l}\cdot\bm{y}+\bm{b}_{l}$, $W_{l}\in\mathbb{R}^{n_{l}\times n_{l-1}}$ and $\bm{b}_{l}\in\mathbb{R}^{n_{l}}$ are the weights and biases of the corresponding layer, respectively, $\bm{x}\in\mathbb{R}^{d}$ is the input, $\bm{\varphi}=(\varphi_{1},\cdots,\varphi_{M})^{T}$ and $\bm{w}=(w^{(1)},\cdots,w^{(M)})^{T}$, $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ is an activation function which can be extended to $\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ as follows: $\sigma(\bm{z})=[\sigma(z_{1}),\cdots,\sigma(z_{n})]^{T}\text{for any}\ \bm{z}=(z_{1},\cdots,z_{n})^{T}\ \in\mathbb{R}^{n}.$ Denote $\theta=\{W_{1},\cdots,W_{L+1},\bm{b}_{1},\cdots,\bm{b}_{L+1}\}\in{\rm\Theta}:=\mathbb{R}^{n_{1}\times n_{0}}\times\cdots\times\mathbb{R}^{n_{L+1}\times n_{L}}\times\mathbb{R}^{n_{1}}\times\cdots\times\mathbb{R}^{n_{L+1}}$ the parameters of the neural network-based functions $\{\varphi_{i}\}_{i=1}^{M}$, and $u(\bm{x};\theta,\bm{w})=\sum\limits_{i=1}^{M}w^{(i)}\varphi_{i}(\bm{x};\theta)$ are the output of neural networks with respect to the input $\bm{x}$ and parameters $\theta$ and $\bm{w}$.
 Next, we will review the general framework of the subspace method based on neural networks for the partial differential equations proposed in [57].
-Firstly, the neural networks architecture introduced above is adopted, and $\theta$ and $w$ are initialized.
-Next, fix the coefficients $w$, and update the subspace basis functions, i.e., update parameters $\theta$ by minimizing a loss function, aiming to obtain basis functions that are better than the initial ones through training.
-After training, calculate the Galerkin projection of the original problem onto the subspace spanned by trained basis functions $\varphi_{1},\cdots,\varphi_{M}$ and obtain an approximate solution to the partial differential equation.
+Firstly, the neural networks architecture introduced above is adopted, and parameters $\theta$ and $\bm{w}$ are initialized.
+Next, train the neural network-based functions, i.e., update parameters $\theta$ by minimizing some loss function, aiming to obtain functions that are better than the initial ones through training.
+After training, find an approximate solution in the subspace spanned by the trained neural network-based functions $\{\varphi_{i}\}_{i=1}^{M}$, i.e., update parameters $\bm{w}$ by solving some linear system, aiming to obtain an approximate solution of the partial differential equation.
 The above process is summarized as a general framework for solving the partial differential equations in [57].
-We also list it here for the completeness, see Algorithm (1) for the details.
-Algorithm 1 General framework of SNN for PDEs
-  1. Initialize the neural networks architecture (4), generate randomly $\theta$, and give $w$.
-  2. Update parameters $\theta$ by minimizing a loss function, and obtain the updated basis functions $\varphi_{1},\cdots,\varphi_{M}$.
-  3. Calculate the Galerkin projection of the PDE onto the subspace $\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$ and obtain an approximate solution.
+We also list it here for the completeness, see Algorithm 1 for the details.
+Algorithm 1 General framework of the subspace method based on neural networks for partial differential equations
+ 1. Initialize the neural network architecture, give and then fix $\bm{w}_{0}$.
+ 2. Define some loss function $\mathcal{L}(\theta,\bm{w}_{0})$ using information contained in the PDEs, then obtain parameters $\theta$ by minimizing the loss function using some optimization method
+$\theta^{\ast}=\mathop{\arg\min}\limits_{\theta\in\rm\Theta}\mathcal{L}(\theta,\bm{w}_{0}),$
+and obtain the updated neural network-based functions $\{\varphi_{i}(\bm{x};\theta^{\ast})\}_{i=1}^{M}$
+ 3.
+Using some classical method to find the approximation for the solution of the original problems in the trained subspace spanned by $\{\varphi_{i}(\bm{x};\theta^{\ast})\}_{i=1}^{M}$, i.e., obtain optimal $\bm{w}^{\ast}$, and then obtain the approximate solution $u(\bm{x};\theta^{\ast},\bm{w}^{\ast})$.
 3 Subspace method based on neural networks for eigenvalue problems
 We consider the following typical eigenvalue problem:
-$\begin{cases}Lu=\lambda u&\text{in }\Omega,\\ u=0&\text{on }\partial\Omega,\end{cases}$
-where the differential operator $L:=-\nabla\cdot(\alpha\nabla)+\beta$ and $\alpha=(\alpha_{ij})_{ij}$ is symmetric positive definite with $\alpha_{ij}\in W^{1,\infty}(\Omega)(i,j=1,\cdots,d)$, and $0\leq\beta\in L^{\infty}(\Omega).$
-The associate weak form of the eigenvalue problem above is as follows: Find $(\lambda,u)\in\mathbb{R}\times H_{0}^{1}(\Omega)$ such that $b(u,u)=1$ and
-$a(u,v)=\lambda b(u,v),\quad\forall v\in H_{0}^{1}(\Omega),$ (5)
+$\begin{cases}Lu=\lambda u&\text{in }\rm\Omega,\\ u=0&\text{on }\partial\rm\Omega,\end{cases}$
+where the differential operator $L:=-\nabla\cdot(\alpha\nabla)+\beta$ and $\alpha=(\alpha_{ij})_{ij}$ is symmetric positive definite with $\alpha_{ij}\in W^{1,\infty}({\rm\Omega})(i,j=1,\cdots,d)$, and $0\leq\beta\in L^{\infty}(\rm\Omega).$
+The associated weak form of the eigenvalue problem above is as follows: Find $(\lambda,u)\in\mathbb{R}\times H_{0}^{1}(\rm\Omega)$ such that $b(u,u)=1$ and
+$a(u,v)=\lambda b(u,v),\quad\forall v\in H_{0}^{1}(\rm\Omega),$ (5)
 which is nothing but (1) with $a(u,v)=(\alpha\nabla u,\nabla v)+(\beta u,v)$ and $b(u,v)=(u,v)$.
-Here, $a(\cdot,\cdot),b(\cdot,\cdot)$ are two positive definite symmetric bilinear forms on $H_{0}^{1}(\Omega)\times H_{0}^{1}(\Omega)$.
+Here, $a(\cdot,\cdot)$ and $b(\cdot,\cdot)$ are two positive definite symmetric bilinear forms on $H_{0}^{1}({\rm\Omega})\times H_{0}^{1}(\rm\Omega)$.
 Remark 1
 It is worth mentioning that our method is also valid for a more general bilinear form $a(\cdot,\cdot)$ that
-$\|v\|_{1}^{2}-C\|v\|_{0}^{2}\lesssim a(v,v),\quad\forall v\in H_{0}^{1}(\Omega)$
+$\|v\|_{1}^{2}-C\|v\|_{0}^{2}\lesssim a(v,v),\quad\forall v\in H_{0}^{1}(\rm\Omega)$
 holds for some constant $C$. (see, e.g., Remark 2.9 in [17])
-3.1 A general framework of SNN for eigenvalue problems
-In this subsection, we extend the subspace method based on neural networks for PDEs, which is proposed in [49] and introduced in Section 2, to eigenvalue problems.
-A neural networks architecture similar to the one introduced in the previous section is adopted.
-The only difference is that the neural network in Section 2 is used to approximate the solution to the boundary value problem, here the neural network is used to approximate the eigenfunctions corresponding to the first $k$ eigenvalues of some partial differential operator.
+Let $V_{h}$ be a finite dimensional subspace of $V$. The finite dimensional discretization of (5) is as follows: Find $(\lambda,u_{h})\in\mathbb{R}\times V_{h}$ such that $b(u_{h},u_{h})=1$ and
+$a(u_{h},v_{h})=\lambda b(u_{h},v_{h}),\quad\forall v_{h}\in V_{h}.$ (6)
+In this paper, we consider to obtain the smallest $k$ eigenvalues and their corresponding eigenfunctions of the eigenvalue problem (5) and discretize the eigenvalue problem (5) in a finite dimensional space spanned by a neural network-based orthogonal basis.
+Our basic idea is as follows: we first use deep learning method to train some neural network-based functions $\{\varphi_{i}\}_{i=1}^{M}$ which is similar to [57]; then, we introduce some dimensionality reduction technique to obtain some orthogonal basis $\{\psi_{i}\}_{i=1}^{K}$ from these trained neural network-based functions $\{\varphi_{i}\}_{i=1}^{M}$; after that, we use some classical method to obtain the approximation solution of the eigenvalue problem (5) in the subspace $V_{h}=\text{span}\{\psi_{1},\cdots,\psi_{K}\}$.
+Usually, there holds $K\ll M$.
+Here, we would like to point out that different from the work in [57], here, we add a step using some dimensionality reduction technique to the trained neural network-based functions.
+In fact, we have done some numerical tests, which show the necessity of using dimensionality reduction technique.
+Some of the results are provided in Appendix B.
+We first give a general framework of our subspace method based on neural networks for solving eigenvalue problems, as in Algorithm 2, and then provide more details for each of the steps in the following parts of this section.
+Here, to make the framework more readable, we combine the training of the neural network-based functions using some deep learning method and obtaining the orthogonal basis from those trained function by some dimensionality reduction technique together and view them as a separate step for obtaining an orthogonal basis.
+Algorithm 2 General framework of subspace method based on neural networks for eigenvalue problems
+ 1. Construct a neural network-based orthogonal basis $\{\psi_{i}\}_{i=1}^{K}$ by some deep learning method and dimensionality reduction technique.
+ 2. Use some classical method, e.g., the Galerkin method, to discretize the eigenvalue problem (5) in the subspace spanned by the basis $\{\psi_{i}\}_{i=1}^{K}$ and obtain an approximate solution.
+3.1 Constructing a neural network-based orthogonal basis
+In this subsection, we introduce how we obtain the neural network-based orthogonal basis $\{\psi_{i}\}_{i=1}^{K}$ by using some deep learning method and dimensionality reduction technique.
+Inspired by the subspace method based on neural networks proposed in [57], we consider to use neural network-based functions to construct basis and train the parameters of those neural network-based functions by minimizing some loss function.
+Furthermore, we observe the difference between solving eigenvalue problems and source problems.
+To solve the algebraic eigenvalue problem resulted from the operator eigenvalue problems with high accuracy, the stiffness matrix and the mass matrix generated by neural network-based basis functions are required to be symmetric and have good condition numbers.
+Usually, the trained neural network-based functions lack good linear independence.
+Consequently, some strategy must be taken to remove the redundant information contained in the trained neural network-based functions.
+In fact, as referred to in the beginning of this section, we have done some tests to show the necessity of using dimensionality reduction technique in Appendix B.
+In addition, to deal with boundary conditions, we also hope the basis satisfies boundary conditions of eigenvalue problem (5) exactly.
+Based on the consideration above, we get the main steps for obtaining a neural network-based orthogonal basis $\{\psi_{i}\}_{i=1}^{K}$ by using some deep learning method and dimensionality reduction technique as follows: First, choose a neural networks architecture with its parameters to be determined; then, modify neural network-based functions to make them satisfy boundary conditions exactly; after that, construct a proper loss function by the neural network-based functions and the information provided by the problem to be solved, and obtain the undetermined parameters in the neural network-based functions by minimizing the loss function; finally, construct a neural network-based orthogonal basis from the trained neural network-based functions by using some dimensionality reduction technique.
+In the following of this subsection, we will introduce the details for each step.
+3.1.1 Neural networks architecture
+The neural networks architecture used here is similar to the one introduced in the previous section.
+The only difference is that the neural networks in Section 2 is used to approximate the solution to the source problem, here the neural networks are used to approximate the eigenfunctions corresponding to the first $k$ eigenvalues of some partial differential operator.
 Therefore, the dimension of the output layer is $k$.
-Thus, the output of the neural network can be presented as follows,
-$\bm{u}(x;\theta,\bm{w})=\begin{pmatrix}u_{1}(x;\theta,w_{1})\\ \cdots\\ u_{k}(x;\theta,w_{k})\end{pmatrix}=\begin{pmatrix}\varphi\cdot w_{1}\\ \cdots\\ \varphi\cdot w_{k}\end{pmatrix}=\begin{pmatrix}\sum\limits_{j=1}^{M}w_{1}^{(j)}\varphi_{j}(x,\theta)\\ \cdots\\ \sum\limits_{j=1}^{M}w_{k}^{(j)}\varphi_{j}(x,\theta)\end{pmatrix}$
-where $\bm{w}=(w_{1},\cdots,w_{k})$, $w_{s}=(w_{s}^{(1)},\cdots,w_{s}^{(M)})^{T}$ denote the coefficients corresponding to the $s$-th output, and $u_{s}(\bm{x};\theta,w_{s})$ denotes the $s$-th output, $s=1,\cdots,k$.
-Following a framework similar to Algorithm 1, we initialize $\theta$ and $w_{1},\cdots,w_{k}$ firstly.
-Then, to handle the boundary condition, appropriate modifications to the basis functions are made according to the specific problem requirements.
-Next, we fix the coefficients $w_{1},\cdots,w_{k}$ and update parameters $\theta$ by minimizing some loss function, from which we obtain the updated basis functions $\varphi_{1},\cdots,\varphi_{M}$.
-Finally, we calculate the Galerkin projection of the original eigenvalue problem onto the subspace spanned by basis functions $\varphi_{1},\cdots,\varphi_{M}$ and obtain an approximate solution to the eigenvalue problem (5).
-The above process is summarized as Algorithm 2.
-Algorithm 2 General framework of SNN for eigenvalue problems
-  1. Initialize the neural networks architecture (4), generate randomly $\theta$, and give $w_{1}$$\cdots$,$w_{k}$, then modify the neural network basis functions according to the boundary conditions.
-  2. Update parameters $\theta$ by minimizing a loss function, and obtain the updated basis functions $\varphi_{1},\cdots,\varphi_{M}$.
-  3. Calculate the Galerkin projection of the eigenvalue problem (5) onto the subspace $\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$ and obtain an approximate solution.
-In the following subsections, we will follow the process outlined in the Algorithm 2 to provide a step-by-step explanation.
-3.2 Initializing and modifying basis functions of the subspace
-3.2.1 Initializing the basis functions
-First, we introduce the initialization of the neural network parameters $\theta$ and the coefficients $w_{1},\cdots,w_{k}$.
-There are various ways to initialize the parameters $\theta$ for initializing the neural network.
-In this paper, we adopt the default initialization strategy provided by PyTorch for neural network parameters $\theta$.
-The coefficients $w_{1},\cdots,w_{k}$ are given and will remain fixed during the training process.
-While there are various methods for assigning the coefficients $w_{1},\cdots,w_{k}$, numerical experiments provided in [49] indicate that the different choices of $w_{1},\cdots,w_{k}$ have no obvious impact on the results.
-Therefore, here, we do not pay much attention on the choice of $w_{1},\cdots,w_{k}$.
+Thus, the output of the neural networks can be presented as follows,
+$\bm{u}(\bm{x};\theta,w)=(\bm{\varphi}\cdot\bm{w}_{1},\cdots,\bm{\varphi}\cdot\bm{w}_{k})^{T}=\left(\sum\limits_{j=1}^{M}w_{1}^{(j)}\varphi_{j}(\bm{x},\theta),\cdots,\sum\limits_{j=1}^{M}w_{k}^{(j)}\varphi_{j}(\bm{x},\theta)\right)^{T},$
+where $w=(\bm{w}_{1},\cdots,\bm{w}_{k})$, $\bm{w}_{s}=(w_{s}^{(1)},\cdots,w_{s}^{(M)})^{T}$ and $u_{s}(\bm{x};\theta,\bm{w}_{s})$ denotes the $s$-th output of the neural networks for $s=1,\cdots,k$.
+Similar to those introduced in Section 2, here, $\bm{\varphi}(\bm{x},\theta)=\sigma\circ\mathcal{T}_{L+1}\circ\cdots\circ\sigma\circ\mathcal{T}_{1}(\bm{x})$, $\sigma$ is an activation function, $\mathcal{T}_{l}(\bm{y})=W_{l}\cdot\bm{y}+\bm{b}_{l}$ and $\theta=\{W_{1},\cdots,W_{L+1},\bm{b}_{1},\cdots,\bm{b}_{L+1}\}$ are the parameters of the neural network-based functions $\{\varphi_{i}\}_{i=1}^{M}$.
+The initialization of the parameters $\theta$ and $\bm{w}_{1},\cdots,\bm{w}_{k}$ are as follows.
+For the parameters $\theta$, we adopt the default initialization strategy provided by PyTorch.
+For the parameters $\bm{w}_{1},\cdots,\bm{w}_{k}$, there are various methods for assigning initial values.
+Numerical experiments provided in [57] indicate that different choices of $\bm{w}_{1},\cdots,\bm{w}_{k}$ have no obvious impact on the results.
+Therefore, here, we do not pay much attention on the choice of $\bm{w}_{1},\cdots,\bm{w}_{k}$.
 We just give some simple choices.
-That is, for a single output, we choose $w=(1,\cdots,1)^{T}\in\mathbb{R}^{M}$.
-For multiple outputs, a random approach is chosen, that is, $w_{1},\cdots,w_{k}\in\mathbb{R}^{M}$ are assigned random values generated from a uniform distribution defined over the interval $[-1,1]$.
-3.2.2 Modifying the basis functions
-For a partial differential equation, some boundary conditions are usually required.
-Since the basis functions constructed by some neural networks are usually not local, it is not so easy to deal with the boundary condition by using the same way as traditional methods.
-Here, we choose to modify the basis functions to make them satisfy the boundary condition.
-For example, for the case of zero boundary condition with $\Omega$ being bounded, we can modify the basis function by the following way.
+That is, for a single output, we choose $\bm{w}=(1,\cdots,1)^{T}\in\mathbb{R}^{M}$.
+For multiple outputs, a random approach is chosen, that is, $\bm{w}_{1},\cdots,\bm{w}_{k}\in\mathbb{R}^{M}$ are assigned random values generated from a uniform distribution defined over the interval [1,1][-1,1].
+3.1.2 Handling the boundary condition
+For eigenvalue problems, some boundary conditions are usually required.
+Since the neural network-based functions are usually nonlocal, it is not so easy to deal with the boundary condition by using the same way as classical methods.
+Here, we choose to modify those functions to make them satisfy the boundary condition.
+For example, for the case of zero boundary condition with $\rm\Omega$ being bounded, we can modify the basis function by the following way.
 We multiply $\varphi_{i}$ by a function $f$ such that
-$\bar{\varphi}_{i}=\varphi_{i}f,\ \bar{\varphi}_{i}(\bm{x})=0,\ \forall\bm{x}\in\partial\Omega,\ \forall i=1,\cdots,M.$
+$\bar{\varphi}_{i}=\varphi_{i}f,\ \bar{\varphi}_{i}(\bm{x})=0,\ \forall\bm{x}\in\partial{\rm\Omega},\ \forall i=1,\cdots,M.$ (7)
 For example, for the case of the 2D unit square domain, we can choose $f(\bm{x})=x_{1}x_{2}(1-x_{1})(1-x_{2})$.
-If $\Omega$ is unbounded, it is necessary to modify the basis functions to ensure that the modified functions are $L^{2}$ integrable.
+If $\rm\Omega$ is unbounded, it is necessary to modify the neural network-based functions to ensure that the modified functions are $L^{2}$ integrable.
 We can achieve this by multiplying $\varphi_{i}$ by a function $f$ such that
-[…]
-Since the Hermite-Gauss quadrature scheme are used to numerically approximate the integrals on unbounded domain in this paper, we can choose $f(\bm{x})=\exp(-\frac{1}{2}\bm{x}^{T}\bm{x})$.
-For simplicity, in the remaining parts of this paper, $\varphi_{i}$ will refer to the modified basis function $\bar{\varphi}_{i}$ as described above.
-3.3 Updating the basis functions
-In this subsection, we will introduce a specific approach to update basis functions by minimizing some loss function $\mathcal{L}$.
-Denote $\bm{w}=(w_{1},\cdots,w_{k})$. For given coefficients $\bm{w}$, parameters $\theta$ can be updated by solving the following minimization problem:
-$\min_{\theta\in\Theta}\mathcal{L}(\theta,\bm{w}).$ (6)
-The gradient descent method is adopted to solve minimization problem (6). That is, we update the parameter by
-$\theta^{(l+1)}=\theta^{(l)}-\eta\nabla_{\theta}\mathcal{L}(\theta^{(l)},\bm{w}),$
-where $\theta^{(l)}$ is the value of $\theta$ in $l$-th epoch and $\eta$ is the learning rate.
-We now introduce our choices for the loss function and the termination criterion used in the above process.
-Before introducing the loss function we used, we first introduce some properties of the eigenvalue problem (5) to give us an idea on how to choose the loss function.
-Denote $V_{k}=\text{span}\{v_{1},\cdots,v_{k}\}$ an arbitrary $k$-dimensional subspace of $V$ and $U_{k}=\text{span}\{u_{1},\cdots,u_{k}\}$ the eigensubspace corresponding to the first $k$ eigenvalues.
-According to the the minimum-maximum principle [6], the sum of the first $k$ eigenvalues and eigenfunctions corresponding to the first $k$ eigenvalues satisfy the following equations,
+$\bar{\varphi}_{i}=\varphi_{i}f,\ \int_{\rm\Omega}\bar{\varphi}_{i}^{2}dx<\infty,\ \forall i=1,\cdots,M.$ (8)
+Since the Hermite-Gauss quadrature scheme is used to numerically approximate the integrals on an unbounded domain in this paper, we can choose $f(\bm{x})=\exp(-\frac{1}{2}\bm{x}^{T}\bm{x})$.
+For simplicity, in the remaining parts of this paper, $\varphi_{i}$ will refer to the modified neural network-based function $\bar{\varphi}_{i}$ as described above.
+3.1.3 The loss function and the training
+In this subsection, we will introduce our choice for the loss function $\mathcal{L}$ and a specific approach to train the neural network-based functions.
+We first introduce our choices for the loss function. Before that, we first introduce some properties of the eigenvalue problem (5) to give us an idea on how to choose the loss function.
+Denote $V_{k}=\text{span}\{v_{1},\cdots,v_{k}\}$ an arbitrary $k$-dimensional subspace of $V$ and $U_{k}=\text{span}\{u_{1},\cdots,u_{k}\}$ the eigensubspace corresponding to the smallest $k$ eigenvalues.
+According to the minimum-maximum principle [6], the sum of the smallest $k$ eigenvalues and eigenfunctions corresponding to the smallest $k$ eigenvalues satisfy the following equations,
 $\sum_{i=1}^{k}\lambda_{i}=\min_{V_{k}=\text{span}\{v_{1},\cdots,v_{k}\}\atop v_{k}\in V,dim(V_{k})=k}\text{Trace}\left(\mathcal{B}^{-1}(v_{1},\cdots,v_{k})\mathcal{A}(v_{1},\dots,v_{k})\right),$ (9)
@@
 Here, $\mathcal{A}$ is the stiffness matrix and $\mathcal{B}$ is the mass matrix, that is, $\mathcal{A}(v_{1},\dots,v_{k})=(a(v_{i},v_{j}))_{ij}\in\mathbb{R}^{k\times k}$, $\mathcal{B}(v_{1},\dots,v_{k})=(b(v_{i},v_{j}))_{ij}\in\mathbb{R}^{k\times k}$.
+We now consider the approximation of the eigenvalue problem (5) on the finite dimensional subspace $V_{h}$ spanned by neural network-based functions.
+Set $V_{h}=\text{span}\{\varphi_{1}(\bm{x};\theta),\cdots,\varphi_{M}(\bm{x};\theta)\}\subset H^{1}_{0}(\rm\Omega)$.
+Set $v_{i}(\bm{x};\theta,\bm{w}_{i})=\sum\limits_{j=1}^{M}w_{i}^{(j)}\varphi_{j}(\bm{x};\theta)$, $i=1,\cdots,k$.
 Motivated by (9) and (10), we then construct the following loss function for finding first $k$ eigenpairs (see also in [53]):
-$\mathcal{L}(\theta,\bm{w})=\text{Trace}(\mathcal{B}^{-1}(v_{1}(x;\theta,w_{1}),\cdots,v_{k}(x;\theta,w_{k}))\mathcal{A}(v_{1}(x;\theta,w_{1}),\cdots,v_{k}(x;\theta,w_{k}))),$
-where $v_{i}(x;\theta,w_{i})=\sum\limits_{j=1}^{M}w_{i}^{(j)}\varphi_{j}(x,\theta)$ is an approximation of the eigenvector corresponding to the $k$-th eigenvalue, $i=1,\cdots,k.$
-As for the choice of termination criteria, we notice that the goal of training is to obtain improved basis functions compared to the initial ones, there is no need to train until the loss is small enough, e.g.
+$\mathcal{L}(\theta,w)=\text{Trace}(\mathcal{B}^{-1}(v_{1}(x;\theta,\bm{w}_{1}),\cdots,v_{k}(x;\theta,\bm{w}_{k}))\mathcal{A}(v_{1}(x;\theta,w_{1}),\cdots,v_{k}(x;\theta,\bm{w}_{k}))).$ (11)
+Now, we turn to introduce how we train the neural network-based functions.
+Similar to Algorithm 1 proposed in [57], for training the neural network-based functions, we first fix the parameter $w$, that is, set $w=w^{(0)}$ with $w^{(0)}$ being given at initialization and then obtain parameters $\theta$ by solving the following minimization problem:
+$\min_{\theta\in\rm\Theta}\mathcal{L}(\theta,w^{(0)}).$ (12)
+The gradient descent method is adopted to solve the minimization problem (12). That is, we update the parameters $\theta$ by
+$\theta^{(l+1)}=\theta^{(l)}-\eta\nabla_{\theta}\mathcal{L}(\theta^{(l)},w^{(0)}),$
+where $\theta^{(l)}$ is the value of $\theta$ in $l$-th epoch and $\eta$ is the learning rate adjusted by the Adam optimizer [28]. Here, when minimizing the loss function, a fixed $w^{(0)}$ is adopted
+We have also briefly explained how the training process affects the accuracy of final solution through some numerical experiments in Appendix A.
+For the choice of termination criteria of the gradient descent method, we notice that the goal of training is to obtain the basis other than the solution, there is no need to train until the loss is small enough, e.g.
 find the $\theta^{\ast}$ s.t.
-$\mathcal{L}(\theta^{\ast},\bm{w})-\min\limits_{\theta\in\Theta}\mathcal{L}(\theta,\bm{w})\leq\epsilon$.
+$\mathcal{L}(\theta^{\ast},w^{(0)})-\min\limits_{\theta\in\rm\Theta}\mathcal{L}(\theta,w^{(0)})\leq\epsilon$.
 Instead, in [57], the following termination criterion is adopted:
-$\frac{\mathcal{L}(\theta,\bm{w})}{\mathcal{L}_{0}(\theta,\bm{w})}\leq\epsilon\ \text{or}\ N_{epoch}\geq N_{max},$
-where $\mathcal{L}_{0}(\theta,\bm{w})$ is the initial loss, $\epsilon$ is the tolerance, $N_{epoch}$ and $N_{max}$ are the current and maximum number of training epochs, respectively.
+$\frac{\mathcal{L}(\theta,w^{(0)})}{\mathcal{L}_{0}(\theta,w^{(0)})}\leq\epsilon\ \text{or}\ N_{epoch}\geq N_{max},$
+where $\mathcal{L}_{0}(\theta,w^{(0)})$ is the initial loss, $\epsilon$ is the tolerance, $N_{epoch}$ and $N_{max}$ are the current and maximum number of training epochs, respectively.
 However, since the minimum value of the loss function we used in this paper is not 0, this criterion may not be suitable for determining when to stop training.
@@
 Therefore, we choose to use the change of the loss during multiple adjacent epochs as termination criterion.
-In this paper, we adopt the change of the loss during 10 adjacent epochs as an indicator.
+In this paper, we adopt the change of the loss during $m$ adjacent epochs as an indicator.
 Any other reasonable choices are also acceptable.
-Let $l_{s}=\mathcal{L}_{s}(\theta,\bm{w})$ be the loss in the $s$-th epoch, $\bar{l}_{s}=\sum\limits_{i=1}^{10}|l_{s-i}$ $-l_{s+1-i}|$ be the change of the loss function during 10 adjacent epochs.
-The the following termination criterion is adopted:
+Let $l_{s}=\mathcal{L}_{s}(\theta,w)$ be the loss in the $s$-th epoch, $\bar{l}_{s}=\sum\limits_{i=1}^{m}|l_{s-i}$ $l_{s+1-i}|$ be the change of the loss function during $m$ adjacent epochs.
+The following termination criterion is adopted:
 $|\frac{\bar{l}_{s}}{l_{s}}|\leq\epsilon,\ \text{or}\ N_{epoch}\geq N_{max},$ (13)
 where $\epsilon$ is a given tolerance, $N_{epoch}$ and $N_{max}$ are the current and maximum training epochs, respectively.
-Under this criterion, the termination of training means that the relative change of the loss function during 10 adjacent epochs has been relatively small, which indicates that the benefits of continued training may be small.
+Under this criterion, the termination of training means that the relative change of the loss function during $m$ adjacent epochs has been relatively small, which indicates that the benefits of continued training may be small.
 Since there is no need to train until the loss is small enough, a relatively big $\epsilon$ and a relatively small $N_{max}$ can be chosen.
-3.4 Galerkin projection of the original eigenvalue problem onto the subspace
-In this subsection, we will introduce the details on how to obtain the finite dimensional approximation of the eigenvalue problem (3) in the subspace $\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$.
-From (3) in Section 2, we know that finding the Galerkin approximation of the eigenvalue problem (5) in subspace $,\text{span}\{\varphi_{1},\cdots$$,\varphi_{M}\}$ can be formulated as the following matrix eigenvalue problem:
-$Aw=\lambda Bw,$ (10)
-where $A=(a(\varphi_{i},\varphi_{j}))_{ij},B=(b(\varphi_{i},\varphi_{j}))_{ij}\in\mathbb{R}^{M\times M},w=(w^{(1)},\cdots,w^{(M)})^{T}\in\mathbb{R}^{M}.$
-To calculate the stiffness matrix and mass matrix, we have to do some numerical integration.
+3.1.4 Dimensionality reduction technique
+In this subsection, we will introduce how we obtain the neural network-based orthogonal basis $\{\psi_{i}\}_{i=1}^{K}$ from the trained neural networks-based functions $\{\varphi_{i}\}_{i=1}^{M}$ which may numerically lack good linear independence while keeping as much necessary information as possible by some dimensionality reduction technique.
+Here, we choose the proper orthogonality decomposition (POD) to achieve this goal.
+We can describe the problem as follows.
+Denote $V=\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$, $V_{k}\subset V$ any $K$-dimensional subspace of $V$, and ${\rm\Pi}_{K}$ the orthogonal projection operator from $V$ to $V_{K}$.
+The problem is to find an optimal $K$-dimensional subspace $V_{K}^{\ast}$ to approximate the original space $V$ good enough.
+There will be two subproblems.
+One is that, for a given dimension of the subspace $K\leq M$, how to find an optimal $K$-dimensional subspace $V_{K}^{\ast}$ to approximate the original space $V$ in a least–squares sense, the other is how to choose a suitable dimension of the subspace $K$.
+For the first problem, i.e.
+we seek an optimal orthogonal projection ${\rm\Pi}_{K}^{\ast}:V\rightarrow V_{K}^{\ast}$ minimizing the projection error of ${\rm\Phi}=(\varphi_{1},\cdots,\varphi_{M})^{T}\in[L^{2}({\rm\Omega})]^{M}$ under the following least–squares distance,
+[…]
+Hence, we need to solve the following minimization problem,
+[…]
+Denote $U=[u_{1},\cdots,u_{K}]\in\mathbb{R}^{M\times K}$ and
+${\rm\Psi}=U^{T}{\rm\Phi}.$ (15)
+We can rewrite the minimization problem (14) to the minimization problem (16) through the following simple derivation,
+[…]
+where $\mathcal{M}=((\varphi_{i},\varphi_{j}))_{i,j}\in\mathbb{R}^{M\times M}$, $\langle x,y\rangle_{\mathcal{M}}=\langle x,\mathcal{M}y\rangle,\|x\|_{\mathcal{M}}=\sqrt{\langle x,x\rangle_{\mathcal{M}}},\forall x,y\in\mathbb{R}^{M}$ and $e_{1},\cdots,e_{M}$ are the standard basis in $\mathbb{R}^{M}$.
+[…]
+Notice that the problem (16) is nothing but the minimization problem considered in the POD method, which has been systematically studied (see [39, 50]), hence we give the following optimal orthogonal projection ${\rm\Pi}_{K}^{\ast}$ of the problem (16) directly,
+$\mathcal{M}^{2}u_{j}=\mu_{j}\mathcal{M}u_{j},j=1,\cdots,K,$ (17)
+[…]
+where $\mu_{1}\geq\cdots\mu_{M}\geq 0$ are the first $M$ eigenvalues of $\mathcal{M}^{2}$ relative to $\mathcal{M}$, $u_{1},\cdots,u_{M}$ are the corresponding eigenvectors.
+The vectors $u_{1},\cdots,u_{K}$ are called POD modes. In practical implementation, we can obtain them by solving the following symmetric $M\times M$ eigenvalue problem
+$\mathcal{M}\bar{u}_{j}=\mu_{j}\bar{u}_{j},j=1,\cdots,K,$
+$u_{j}=\frac{1}{\sqrt{\mu_{j}}}\bar{u}_{j},j=1,\cdots,K.$
+Remark 2
+In actual implementation, the condition number of $\mathcal{M}$ is usually very large, and it is difficult to obtain a high-precision solution by directly solving the eigenvalue problem.
+Instead, we choose to calculate the singular value of $\mathcal{M}^{\frac{1}{2}}$, from which we can also get $\mu_{j},u_{j},j=1,\cdots,K.$ By the way, we only need to deal with a problem whose condition number becomes the square root of the original problem.
+In addition, we do not need to pay any extra cost for calculating $\mathcal{M}^{\frac{1}{2}}$, because it has a natural approximation as follows.
+Denote $\Phi_{h}=(\varphi_{i}(\bm{x}^{(j)}))_{ij}\in\mathbb{R}^{M\times n},W_{h}=diag(\rho_{1},\cdots,\rho_{n})\in\mathbb{R}^{n\times n}$, where $\{\bm{x}^{(i)},\rho_{i}\}_{i=1}^{n}$ represents the numerical integration point and weight, respectively.
+For $\mathcal{M}$, we have the following approximation, $\mathcal{M}\approx\Phi_{h}W_{h}\Phi_{h}^{T}$.
+Hence, we can use $W_{h}^{\frac{1}{2}}\Phi_{h}^{T}$ to approximate $\mathcal{M}^{\frac{1}{2}}$.
+Regarding the choice of the subspace dimension $K$, since (18) shows the approximate error of the optimal $K$-dimensional subspace is related to the sum of the last MKM-K eigenvalues of the matrix $\mathcal{M}$ (or the square root of the eigenvalues), it’s reasonable to choose a small enough $K$ such that the first $K$ eigenvalues (or the square root of the eigenvalues) account for a high proportion in the overall, i.e.,
+$K=\mathop{\arg\min}\{I(N):I(N)\geq 1-\gamma\},$
+where $I(N)=\frac{\sum\limits_{j=1}^{N}\mu_{j}}{\sum\limits_{j=1}^{M}\mu_{j}}$ or $I(N)=\frac{\sum\limits_{j=1}^{N}\sqrt{\mu_{j}}}{\sum\limits_{j=1}^{M}\sqrt{\mu_{j}}}$ and $\gamma$ is a given parameter.
+In our experiment, we find that the indicator $I(N)$ defined by the latter works better.
+To see the necessity of employing dimensionality reduction technique, we have done some numerical tests, whose results are present in Appendix B.
+3.1.5 Algorithm
+By summarizing the steps above, we obtain our algorithm for constructing the neural network-based orthogonal basis by deep learning method and dimensionality reduction technique, see Algorithm 3 for the details.
+Algorithm 3 Constructing the neural network-based orthogonal basis
+0: the neural network structure: $L,n_{0},n_{1},\cdots,n_{L},M$; parameters used for training the basis: $\epsilon,m,N_{\max},\gamma$; number of eigenpairs needed: $k$
+0:  the number of orthogonal basis functions $K$ and the basis $\{\psi_{i}(\bm{x},\theta^{\ast})\}_{i=1}^{K}.$
+ 1. Choose a neural network architecture and modify neural network-based functions to make it satisfy the boundary condition.
+ 2. Train the neural network-based functions $\{\varphi_{i}(\bm{x},\theta)\}_{i=1}^{M}$ by minimizing the loss function (11) until (13) is satisfied.
+ 3.
+Obtain the neural network-based orthogonal basis $\{\psi_{i}(\bm{x},\theta^{\ast})\}_{i=1}^{K}$ from the trained neural network-based functions $\{\varphi_{i}(\bm{x},\theta^{\ast})\}_{i=1}^{M}$ by POD method.
+3.2 Galerkin discretization in the neural network-based subspace
+In this subsection, we will introduce the details on how we discretize the eigenvalue problem (5) in the subspace spanned by the neural network-based orthogonal basis $\{\psi_{i}(\bm{x},\theta^{\ast})\}_{i=1}^{K}$.
+Similar to the introduction in Section 2, we consider the following finite dimensional discretization of (5): Find $(\lambda,u_{h})\in\mathbb{R}\times V_{h}$ such that $b(u_{h},u_{h})=1$ and
+$a(u_{h},v_{h})=\lambda b(u_{h},v_{h}),\quad\forall v_{h}\in V_{h}.$ (19)
+Let $V_{h}=\text{span}\{\psi_{1}(\bm{x};\theta^{\ast}),\cdots,\psi_{K}(\bm{x};\theta^{\ast})\}$.
+Set $u_{NN}(\bm{x};\theta^{\ast},\bm{w})=\sum\limits_{j=1}^{K}w^{(j)}\psi_{j}(\bm{x};\theta^{\ast})$, and substitute $u_{NN}(\bm{x};\theta^{\ast},\bm{w})$ into (19) and choose $v_{h}(\bm{x})=\psi_{i}(\bm{x};\theta^{\ast})$, we get
+$\sum\limits_{j=1}^{K}w^{(j)}a(\psi_{j},\psi_{i})=\lambda\sum\limits_{j=1}^{K}w^{(j)}b(\psi_{j},\psi_{i}),i=1,\cdots,K.$ (20)
+If we set $A=(a(\psi_{i}(\bm{x};\theta^{\ast}),\psi_{j}(\bm{x};\theta^{\ast})))_{ij},B=((\psi_{i}(\bm{x};\theta^{\ast}),\psi_{j}(\bm{x};\theta^{\ast})))_{ij}\in\mathbb{R}^{K\times K},$ $\bm{w}=(w^{(1)},\cdots,w^{(K)})^{T}\in\mathbb{R}^{K}$, then (20) can be written as the following algebraic eigenvalue problem
+$A\bm{w}=\lambda B\bm{w}.$ (21)
+To calculate the stiffness matrix $A$ and the mass matrix $B$, we have to do some numerical integration.
 For simplicity, we take $a(u,v)=(\nabla u,\nabla v),b(u,v)=(u,v)$ as an example to show how to compute each element of the stiffness matrix and mass matrix by numerical integration.
-They are calculated as follows:
-$(\varphi_{i},\varphi_{j})\approx\sum\limits_{l=1}^{n}\rho_{l}\varphi_{i}(\bm{x}^{(l)})\varphi_{j}(\bm{x}^{(l)}),$ (11)
-$(\nabla\varphi_{i},\nabla\varphi_{j})\approx\sum\limits_{l=1}^{n}\rho_{l}\nabla\varphi_{i}(\bm{x}^{(l)})\nabla\varphi_{j}(\bm{x}^{(l)}).$ (12)
+Since $\theta^{\ast}$ is fixed, for simplicity, we will omit $\theta^{\ast}$ and denote $\psi_{i}(\bm{x},\theta^{\ast})$ as $\psi_{i}(\bm{x})$.They are calculated as follows:
+$(\psi_{i},\psi_{j})\approx\sum\limits_{l=1}^{n}\rho_{l}\psi_{i}(\bm{x}^{(l)})\psi_{j}(\bm{x}^{(l)}),(\nabla\psi_{i},\nabla\psi_{j})\approx\sum\limits_{l=1}^{n}\rho_{l}\nabla\psi_{i}(\bm{x}^{(l)})\nabla\psi_{j}(\bm{x}^{(l)}).$
 Here, $\{\bm{x}^{(l)}\}_{l=1}^{n}$ are the numerical integration points, $\{\rho_{l}\}_{l=1}^{n}$ are the weights for each numerical integration point and $n$ is the number of numerical integration points.
-The first $k$ eigenpairs $(\lambda_{h,1},w_{1}),\cdots,(\lambda_{h,k},w_{k})$ can be obtained by solving the matrix eigenvalue problem (10) using some eigensolver [8, 37], then the approximation of the first $k$ eigenpairs $(\lambda_{h,1},u_{h,1}),\cdots,$ $(\lambda_{h,k},u_{h,k})$ are obtained by setting $u_{h,i}=\sum\limits_{j=1}^{M}w_{i}^{(j)}\varphi_{j}(x,\theta),i=1,\cdots,k.$
-3.5 Limitations
-Unfortunately, some simple tests show that simply applying Algorithm 2 to solution of some eigenvalue problems can not obtained results as accurate as solution of some source problems.
-In fact, we have applied the Algorithm 2 to the 2D Laplace eigenvalue problem on a square domain $[0,1]^{2}$ as introduced in Section 6, aiming to solve the first 15 eigenpairs.
-The neural networks architecture is constructed by a fully connected neural network with 3 hidden layers and a subspace layer with each hidden layer containing 100 neurons and subspace layer containing $M$ neurons.
-We choose $\epsilon=10^{-3}$ and $N_{max}=5000$ in the termination criterion and adopt Legendre-Gauss integration scheme with 32 integration points in each dimension.
-For solving the matrix eigenvalue problem (10), we adopt the solver $scipy.linalg.eig$ in Python.
-Numerical experiments for different $M$ have been carried and the results are presented in Table 1.
-We denote $err_{\lambda,l}$ the relative error the $l$-th eigenvalue, which is defined by $err_{\lambda,l}:=\frac{|\lambda_{l}-\lambda_{h,l}|}{|\lambda_{l}|}$, where $\lambda_{l}$ is the reference eigenvalue.
-From Table 1 we can see that even when the dimension of the subspace $M$ is small, our method can still obtain approximate eigenvalues with high accuracy.
-However, as the increase of $M$, the accuracy of eigenvalues is not improved but rapidly deteriorates, contrary to our expectations.
-Table 1: The relative errors of the first 15 eigenvalues obtained by Algorithm 2 with different $M$ for the 2D Laplace eigenvalue problem on a square domain $[0,1]^{2}$.
+The smallest $k$ eigenvalues and their corresponding eigenfunctions $(\lambda_{h,1},\bm{w}_{1}),\cdots,$ $(\lambda_{h,k},\bm{w}_{k})$ can be obtained by solving the algebraic eigenvalue problem (21) using some eigensolver [9, 42], then the approximation of the $k$ eigenpairs $(\lambda_{h,1},u_{h,1}),\cdots,$ $(\lambda_{h,k},u_{h,k})$ are obtained by setting $u_{h,i}=\sum\limits_{j=1}^{M}w_{i}^{(j)}\psi_{j}(x,\theta^{\ast}),i=1,\cdots,k.$
+4 Numerical experiments
+In this section, we will use four typical eigenvalue problems, the Laplace eigenvalue problem, the decoupled harmonic oscillator, the coupled harmonic oscillator and the hydrogen atom to demonstrate the performance of the methods proposed in our paper.
+The algorithms are implemented based on the deep learning framework PyTorch.
+All the numerical experiments are carried out on the LSSC-IV platform at the Academy of Mathematics and Systems Science, Chinese Academy of Sciences.
+The following parameter settings are used in our numerical experiments: The number of eigenpairs $k$ is set to 15 in all examples.
+The neural networks architecture is implemented by a fully connected neural network with 3 hidden layers and a subspace layer with each hidden layer containing 100 neurons and the subspace layer containing $M$ neurons.
+We choose the activation function $\sigma(x)=\sin(x).$ During the initialization part, the default initialization strategy is adopted to initialize parameters $\theta$ and the random values generated according to a uniform distribution in the range [-1, 1] are applied to initialize parameters $\bm{w}_{1},\cdots,\bm{w}_{k}$.
+The Adam optimizer with the default learning rate is used for training.
+In all experiments, we adopt the termination criteria introduced in (13) with the parameter $N_{max}=5000$ and $m=10$.
+Different $\epsilon$ are chosen for different problems to achieve appropriate training (i.e., fewer epochs while maintaining accuracy).
+The solver $scipy.linalg.eigh$ in Python is used for solving the algebraic eigenvalue problem (21).
+For reproducibility, all the random seeds used in the numerical experiment are set to 1.
+For simplicity, we denote our subspace method based on neural networks for eigenvalue problems, Algorithm 2, as SNN-EP.
+To show the advantage of our new method to existing neural network-based methods, we compare our method with other neural network-based method, the one proposed in [53], which can be viewed as a variant of DRM capable of solving multiple eigenpairs simultaneously, and denote it as DRM-M.
+For the fairness of the comparison, we set the same neural networks architecture, initialization method, and the optimizer.
+The epochs in SNN-EP are determined by the termination criterion (13), while those in DRM-M are set to $25000$.
+Setting the epochs big enough to do the training is a common usage when using deep learning method to solve PDEs.
+The relative error of the approximate eigenvalues and eigenfunctions are defined as follows:
+$err_{\lambda,l}:=\frac{|\lambda_{l}-\lambda_{h,l}|}{|\lambda_{l}|},err_{L^{2},l}:=\frac{\|u_{l}-u_{h,l}\|_{L^{2}({\rm\Omega})}}{\|u_{l}\|_{L^{2}({\rm\Omega})}},err_{H^{1},l}:=\frac{\|u_{l}-u_{h,l}\|_{H^{1}({\rm\Omega)}}}{\|u_{l}\|_{H^{1}({\rm\Omega})}},$
+where $\lambda_{l}$ and $u_{l}$ are reference eigenvalues and eigenfunctions, $l=1,\cdots,k$.
+4.1 Laplace eigenvalue problem
+We first consider the following 2D Laplace eigenvalue problem on a square domain: Find $(\lambda,u)\in\mathbb{R}\times H_{0}^{1}(\rm\Omega)$ such that $\int_{\rm\Omega}|u|^{2}d\rm\Omega=1$ and
+$\begin{cases}-{\rm\Delta}u=\lambda u&\text{in }\rm\Omega,\\ u=0&\text{on }\partial\rm\Omega,\end{cases}$
+where ${\rm\Omega}=[0,1]^{2}$. The exact eigenvalues are as follows:
+$\lambda_{n_{1},n_{2}}=\pi(n_{1}^{2}+n_{2}^{2}),n_{1},n_{2}=1,2,\cdots,$
+and the corresponding eigenfunctions are
+$u_{n_{1},n_{2}}(x_{1},x_{2})=\sqrt{2}\sin\left(\pi n_{1}x_{1}\right)\sin\left(\pi n_{2}x_{2}\right),n_{1},n_{2}=1,2,\cdots.$
+For using SNN-EP to solve the above 2D Laplace eigenvalue problem, since $\rm\Omega$ is bounded, we choose the Legendre-Gauss integration scheme with 32 integration points in each dimension.
+For training, we set $\epsilon=10^{-3}$.
+And for the dimensionality reduction part, we set $\gamma=10^{-11}$.
+The dimension of the reduced subspace and the relative errors of the smallest 15 eigenvalues obtained by SNN-EP with different $M$ are presented in Figure 1.
+It can be observed that when $M$ is small, the dimension of the reduced subspace $K$ increases significantly with the increase of $M$.
+At this stage, the relative errors of the smallest 15 eigenvalues rapidly decay to below $10^{-12}$.
+When $M$ reaches $300$, the increase of $K$ slows down significantly and the relative errors of the smallest 15 eigenvalues remain almost unchanged.
+Therefore, we only show results obtained by setting $M=300$ for both SNN-EP and DRM-M.
+Figure 1: The dimension of the reduced subspace and the relative errors of the smallest 15 eigenvalues for 2D Laplace eigenvalue problem obtained by SNN-EP with different $M$.
+The relative errors of the smallest $15$ eigenpairs obtained using SNN-EP and DRM-M with $M=300$ are presented in Table 1, and epochs required for the training process of SNN-EP are shown in Table 2.
+It can be seen that the accuracy obtained by DRM-M reaches only $10^{-8}$ for eigenvalues and $10^{-4}$ for eigenfunctions after $25000$ epochs, while our method SNN-EP achieves accuracy below $10^{-12}$ for eigenvalues and nearly $10^{-7}$ for eigenfunctions, only using $100$ to $400$ epochs.
+Table 1: The relative errors of the smallest 15 eigenvalues for the 2D Laplace eigenvalue problem obtained by setting $M=300$.
+SNN-EP DRM-M
+$k$ ($n_{1},n_{2}$) $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$ $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$
+1 (1,1) 3.960e-15 2.799e-08 3.912e-07 3.868e-07 1.222e-04 7.086e-04
+2 (2,1) 4.996e-14 4.711e-08 4.239e-07 1.325e-07 1.135e-04 4.374e-04
+3 (1,2) 1.526e-13 5.802e-08 5.366e-07 2.186e-07 1.249e-04 6.110e-04
+4 (2,2) 1.449e-13 6.954e-08 5.151e-07 7.583e-08 9.439e-05 3.828e-04
+5 (3,1) 1.523e-13 8.539e-08 5.753e-07 1.126e-07 2.027e-04 5.936e-04
+6 (1,3) 1.892e-13 9.419e-08 6.128e-07 1.608e-07 1.709e-04 4.993e-04
+7 (3,2) 1.256e-13 1.137e-07 6.993e-07 8.341e-08 1.899e-04 4.586e-04
+8 (2,3) 3.792e-13 1.284e-07 7.331e-07 1.163e-07 1.825e-04 4.520e-04
+9 (4,1) 9.622e-14 1.416e-07 7.757e-07 3.046e-07 1.360e-03 1.962e-03
+10 (1,4) 2.531e-13 1.436e-07 7.976e-07 1.193e-06 5.145e-04 9.427e-04
+11 (3,3) 1.923e-13 1.774e-07 9.253e-07 1.106e-07 2.519e-04 4.826e-04
+12 (4,2) 1.518e-13 1.737e-07 8.788e-07 2.862e-07 1.372e-03 1.908e-03
+13 (2,4) 5.876e-13 2.258e-07 1.181e-06 1.039e-06 5.832e-04 9.282e-04
+14 (4,3) 2.919e-13 2.558e-07 1.096e-06 3.565e-07 1.184e-03 1.516e-03
+15 (3,4) 4.750e-13 3.231e-07 1.435e-06 6.469e-07 8.103e-04 1.084e-03
+Table 2: The epochs needed for using SNN-EP with different $M$ to solve the 2D Laplace eigenvalue problem.
+M 50 100 150 200 250 300 350 400 450 500 550 600
+Epoch 177 195 190 342 160 165 171 160 271 168 200 191
+4.2 Harmonic oscillator problems
+In this subsection, we study the following 2D harmonic oscillator: Find $(\lambda,u)\in\mathbb{R}\times H_{0}^{1}(\rm\Omega)$ such that $\int_{\rm\Omega}|u|^{2}d\rm\Omega=1$ and
+$\frac{1}{2}{\rm\Delta}u+Vu=\lambda u\quad\text{in}\quad\rm\Omega,$
+where ${\rm\Omega}=\mathbb{R}^{2},V(x_{1},x_{2})=\frac{1}{2}(a_{11}x_{1}^{2}+2a_{12}x_{1}x_{2}+x_{22}^{2})$.
+Two concrete cases are considered here.
+For the first case, we use the decoupled harmonic oscillator potential which is most common used, i.e., $a_{11}=a_{22}=1,a_{12}=0$.
+For the second case, we adopt a coupled harmonic oscillator potential which is used in [32] and maintain the same accuracy of the coefficients $a_{11},a_{12},a_{22}$ as [32], i.e.
+$a_{11}=0.8851,a_{12}=-0.1382,a_{22}=1.1933$.
+Case 1: decoupled case. The exact eigenvalues are given as follows:
+$\lambda_{n_{1},n_{2}}=(\frac{1}{2}+n_{1})+(\frac{1}{2}+n_{2}),n_{1},n_{2}=0,1,\cdots,$
+and the corresponding eigenfunctions are
+$u_{n_{1},n_{2}}(x_{1},x_{2})=\mathcal{H}_{n_{1}}(x_{1})e^{-x_{1}^{2}/2}\mathcal{H}_{n_{2}}(x_{2})e^{-x_{2}^{2}/2},n_{1},n_{2}=0,1,\cdots,$
+where $\mathcal{H}_{n}$ is Hermite polynomials [5].
+For using SNN-EP to solve the 2D decoupled harmonic oscillator above, since $\rm\Omega$ is $\mathbb{R}^{2}$, the Hermite-Gauss integration scheme is chosen to compute the unbounded integral, with $99$ integration points in each dimension.
+In the training part, we take $\epsilon=10^{-2}$.
+In the dimensionality reduction part, we take $\gamma=10^{-10}$.
+The dimension of the reduced subspace and the relative errors of the smallest 15 eigenvalues obtained by using different $M$ are presented in Figure 2.
+It can be observed that when $M$ is small, the dimension of the reduced subspace $K$ increases significantly with the increase of $M$.
+At this stage, the relative errors of the smallest $15$ eigenvalues rapidly decay to below $10^{-11}$.
+When $M$ reaches $900$, the increase of the dimension of the reduced subspace $K$ slows down significantly and the relative errors of the smallest $15$ eigenvalues remain almost unchanged.
+Therefore, we only show results obtained by setting $M=900$ for both SNN-EP and DRM-M.
+Figure 2: The dimension of the reduced subspace and the relative errors for the 2D decoupled harmonic oscillator obtained by SNN-EP with different $M$.
+The relative errors of the smallest $15$ eigenpairs obtained using SNN-EP and DRM-M with $M=900$ are presented in Table 3.
+Table 4 provides the number of epochs needed for our method SNN-EP with various $M$.
+Here, it is evident that the accuracy obtained by SNN-EP is significantly superior to that obtained by DRM-M, while SNN-EP requires far fewer epochs than DRM-M.
+Furthermore, our results are also superior to the existing results in literature.
+For example, an accuracy of $10^{-2}$ is achieved using machine learning methods based on Monte Carlo integration, as demonstrated in [32].
+Additionally, our accuracy surpasses that of [53], which employs numerical integration based on TNN to achieve high-precision calculations.
+Compared to [53], which requires $500,000$ epochs, our method achieves the desired accuracy with only $20$ epochs.
+Table 3: The relative errors of the smallest $15$ eigenvalues of the 2D decoupled harmonic oscillator obtained by setting $M=900$.
+SNN-EP DRM-M
+$k$ ($n_{1},n_{2}$) $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$ $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$
+1 (0,0) 7.294e-14 3.977e-08 2.373e-07 2.449e-09 2.062e-05 3.913e-05
+2 (0,1) 8.005e-14 6.304e-08 3.012e-07 4.052e-09 3.637e-05 6.912e-05
+3 (1,0) 8.926e-14 6.095e-08 2.963e-07 8.495e-09 5.009e-05 9.015e-05
+4 (0,2) 3.449e-14 1.184e-07 4.473e-07 3.324e-09 4.887e-05 7.840e-05
+5 (1,1) 4.737e-14 9.585e-08 3.648e-07 4.008e-09 5.239e-05 8.106e-05
+6 (2,0) 1.198e-13 9.863e-08 4.040e-07 6.147e-09 4.121e-05 6.859e-05
+7 (0,3) 7.283e-14 3.044e-07 1.035e-06 3.407e-09 7.011e-05 1.035e-04
+8 (1,2) 9.104e-14 1.458e-07 5.599e-07 3.609e-09 6.003e-05 8.251e-05
+9 (2,1) 2.474e-13 1.632e-07 5.801e-07 7.057e-09 6.483e-05 9.069e-05
+10 (3,0) 7.412e-13 2.373e-07 8.030e-07 8.040e-09 7.635e-05 1.054e-04
+11 (1,3) 3.457e-12 1.312e-06 3.803e-06 6.594e-09 1.537e-04 1.734e-04
+12 (2,2) 4.765e-12 9.109e-07 2.550e-06 7.201e-09 1.784e-04 2.085e-04
+13 (3,1) 5.116e-12 1.221e-06 3.586e-06 1.029e-08 1.329e-04 1.632e-04
+14 (0,4) 1.962e-11 1.923e-06 5.044e-06 1.426e-08 1.176e-04 1.424e-04
+15 (4,0) 2.506e-11 1.185e-06 3.308e-06 2.001e-08 1.511e-04 1.772e-04
+Table 4: The epochs needed for using SNN-EP with different $M$ to solve the 2D decoupled harmonic oscillator.
+M 100 200 300 400 500 600 700 800 900 1200 1500 1800
+Epoch 20 21 23 23 22 20 20 21 21 21 20 22
+Case 2: coupled case. The coupled harmonic oscillator defined by the above potential energy term can be decoupled into the following diagonal form:
+$V(y_{1},y_{2})=\frac{1}{2}(\mu_{1}y_{1}^{2}+\mu_{2}y_{2}^{2}),$
+where $y_{1}=-0.9339352418x_{1}-0.3574422527x_{2}$ and $y_{2}=0.3574422527x_{1}-0.93393524$ $18x_{2}$, $\mu_{1}=0.8322071257$, $\mu_{2}=1.2461928742$.
+Based on the decoupled form, the exact eigenvalues are given as follows:
+$\lambda_{n_{1},n_{2}}=(\frac{1}{2}+n_{1})\mu_{1}^{1/2}+(\frac{1}{2}+n_{2})\mu_{2}^{1/2},n_{1},n_{2}=0,1,\cdots,$
+and the corresponding eigenfunctions are
+$u_{n_{1},n_{2}}(x_{1},x_{2})=\mathcal{H}_{n_{1}}(\mu_{1}^{1/4}y_{1})e^{-\mu_{1}^{1/2}y_{1}^{2}/2}\mathcal{H}_{n_{2}}(\mu_{2}^{1/4}y_{2})e^{-\mu_{2}^{1/2}y_{2}^{2}/2},n_{1},n_{2}=0,1,\cdots.$
+For using SNN-EP to solve the 2D coupled harmonic oscillator above, since $\rm\Omega$ is $\mathbb{R}^{2}$, we choose the Hermite-Gauss integration scheme, with $99$ integration points in each dimension.
+In the training part, we take $\epsilon=10^{-2}$.
+In the dimensionality reduction part, we take $\gamma=10^{-10}$.
+The dimension of the reduced subspace and the relative errors of the smallest $15$ eigenvalues obtained by SNN-EP with different $M$ are presented in Figure 3.
+A numerical performance similar to the decoupled harmonic oscillator can be observed.
+Accordingly, we present detailed results for $M=900$ for both SNN-EP and DRM-M.
+Figure 3: The dimension of the reduced subspace and the relative errors of eigenvalues of the 2D coupled harmonic oscillator obtained by SNN-EP with different $M$.
+The relative errors of the smallest $15$ eigenpairs of the 2D coupled harmonic oscillator obtained by using SNN-EP and DRM-M with $M=900$ are shown in Table 5.
+Similar to the decoupled harmonic oscillator, our method obtains approximate solutions with higher accuracy while requiring fewer epochs during the training, as shown in Table 6, compared to DRM-M and other existing results in [32] and [53].
+Table 5: The relative errors of the smallest $15$ eigenvalues of the 2D coupled harmonic oscillator obtained by setting $M=900$.
+SNN-EP DRM-M
+$k$ ($n_{1},n_{2}$) $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$ $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$
+1 (0,0) 2.330e-11 4.430e-08 2.520e-07 1.154e-07 1.361e-04 2.723e-04
+2 (1,0) 1.979e-11 6.415e-08 3.183e-07 2.964e-08 9.442e-05 1.736e-04
+3 (0,1) 2.647e-11 7.323e-08 3.517e-07 2.188e-08 9.079e-05 1.600e-04
+4 (2,0) 1.939e-11 2.710e-07 1.011e-06 6.870e-08 1.912e-04 2.966e-04
+5 (1,1) 2.348e-11 1.704e-07 6.191e-07 1.830e-08 9.054e-05 1.414e-04
+6 (0,2) 2.740e-11 1.188e-07 4.584e-07 1.826e-08 1.172e-04 1.754e-04
+7 (3,0) 2.829e-11 1.107e-06 3.355e-06 5.469e-08 1.966e-04 2.674e-04
+8 (2,1) 2.312e-11 4.443e-07 1.350e-06 4.192e-08 2.110e-04 2.704e-04
+9 (1,2) 2.545e-11 2.514e-07 8.112e-07 1.655e-08 1.173e-04 1.616e-04
+10 (0,3) 2.798e-11 1.478e-07 5.192e-07 1.759e-08 1.341e-04 1.743e-04
+11 (4,0) 2.524e-10 6.010e-06 1.645e-05 9.014e-08 3.495e-04 4.336e-04
+12 (3,1) 1.159e-10 4.032e-06 1.036e-05 6.991e-08 3.703e-04 4.364e-04
+13 (2,2) 4.021e-11 1.626e-06 4.390e-06 3.048e-08 2.252e-04 2.719e-04
+14 (1,3) 3.324e-11 1.120e-06 2.951e-06 4.191e-08 2.746e-04 3.394e-04
+15 (0,4) 2.874e-11 3.594e-07 1.063e-06 3.972e-08 2.894e-04 3.600e-04
+Table 6: The epochs needed for using SNN-EP with different $M$ to deal with the 2D coupled harmonic oscillator.
+M 100 200 300 400 500 600 700 800 900 1200 1500 1800
+Epoch 20 21 22 23 22 20 20 21 20 21 20 23
+4.3 Hydrogen atom
+We consider the following Schrodinger equation for Hydrogen atom in 2D: Find $(\lambda,u)\in\mathbb{R}\times H_{0}^{1}(\rm\Omega)$ such that $\int_{\rm\Omega}|u|^{2}d\rm\Omega=1$ and
+$\frac{1}{2}{\rm\Delta}u+Vu=\lambda u\quad\text{in}\quad\rm\Omega,$
+where ${\rm\Omega}=\mathbb{R}^{2},V(x_{1},x_{2})=-\frac{1}{\sqrt{x_{1}^{2}+x_{2}^{2}}}$.
+The exact eigenvalues are as follows:
+$\lambda_{n_{1},n_{2}}=-\frac{2}{(2|n_{1}|+2n_{2}+1)^{2}},|n_{1}|,n_{2}=0,1,\cdots,,$
+and the corresponding eigenfunctions in spherical coordinates are
+$u_{n_{1},n_{2}}(r,\theta)=e^{in_{1}\theta}e^{-\frac{2}{2|n_{1}|+2n_{2}+1}r}r^{|n_{1}|}\mathcal{P}_{|n_{1}|,n_{2}}(r),|n_{1}|,n_{2}=0,1,\cdots,$
+where $\mathcal{P}_{m,n}(r)=\sum\limits_{i=0}^{n}a_{i}r^{i},$ $a_{i+1}=\frac{\beta(2m+2i+1)-2}{(i+1)(i+2m+1)}a_{i}$, $\beta=\frac{2}{2|m|+2n+1},i=0,1,\cdots,n-1$ and $a_{0}$ is chosen to normalize $u_{n_{1},n_{2}}(r,\theta)$.
+To deal with the singularity of the potential function $V$, we generate numerical integration points in spherical coordinates, i.e., $\int_{\mathbb{R}^{2}}g(x_{1},x_{2})dx_{1}dx_{2}=\int_{0}^{2\pi}\int_{0}^{+\infty}g(r\cos\theta,$ $r\sin\theta)r^{2}sin\theta drd\theta,$ where $g$ is a general function.
+For the integration along $r$, we choose the Hermite-Gauss integration scheme, with 99 integration points.
+For the integration along $\theta$, we choose the Legendre-Gauss integration scheme, with 4 integration points in each subinterval and the number of subintervals $20$ .
+In training part, we take $\epsilon=10^{-2}$.
+In the dimensionality reduction part, we take $\gamma=10^{-10}$.
+Different from the previous examples, here, we apply 5 different modification functions $f_{0},\cdots,f_{4}$ to each neural network-based function $\varphi_{j}$.
+Then we train the parameters and perform dimensionality reduction operation to the modified functions $\{f_{i}\varphi_{j}\}_{i=0,j=1}^{4,M}$, here $f_{i}(x_{1},x_{2})=e^{-\frac{2}{2i+1}{\sqrt{x_{1}^{2}+x_{2}^{2}}}},i=0,1,2,3.$
+The dimension of the reduced subspace and the relative errors of the smallest 15 eigenvalues obtained by using different $M$ are presented in Figure 4.
+It can be observed that when $M$ is small, the dimension of the reduced subspace $K$ increases significantly with the increase of $M$.
+At this stage, the relative errors of the smallest $15$ eigenvalues rapidly decay to below $10^{-12}$.
+When $M$ reaches $600$, the increase of the dimension of the reduced subspace $K$ slows down significantly.
+When $M$ reaches $1000$, the relative errors of the smallest $15$ eigenvalues remain almost unchanged.
+Therefore, we only show results obtained by setting $M=1000$ for both SNN-EP and DRM-M in Table 7.
+From these results, we can see clearly that the accuracy obtained by SNN-EP is significantly superior to that obtained by DRM-M, while SNN-EP requires much fewer epochs as listed in Table 8.
+It should be pointed out that the selection of the above $f_{i}$ is too specific.
+In fact, we can also choose $f_{i}$ in the following more general way to obtain $\{f_{i}\varphi_{j}\}_{i=1,j=1}^{N,M}$, e.g.
+$f_{i}(\bm{x})=e^{-\beta_{i}{\|x\|}}$, $i=1,\cdots,N.$ Here, $\beta_{i}$ is randomly chosen from a uniform distribution and $N$ represents the number of randomly selected $\beta_{i}$.
+For the randomly selected $\beta_{i}$, we have also done some numerical experiments, whose results are presented in Figure 5.
+These results show that as long as we use enough number of $\beta_{i}$, even random ones can achieve a similar effect, and for this problem, NN=10 is sufficient.
+Figure 4: The dimension of the reduced subspace and the relative error of eigenvalues of the Schrodinger equation for Hydrogen atom obtained by SNN-EP with different $M$.
+Table 7: The relative errors of the smallest $15$ eigenvalues of the Schrodinger equation for Hydrogen atom obtained by SNN-EP with $M=1000$.
+SNN-EP DRM-M
+$k$ ($n_{1},n_{2}$) $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$ $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$
+1 ( 0,0) 2.770e-12 6.092e-09 1.642e-08 2.847e-07 1.144e-04 1.370e-04
+2 ( 0,1) 1.392e-13 5.620e-08 3.046e-07 6.322e-04 1.047e-02 5.763e-02
+3 ( 1,0) 2.244e-13 6.228e-08 8.909e-08 1.479e-05 2.991e-03 6.391e-03
+4 (-1,0) 1.101e-12 4.646e-08 5.109e-08 6.761e-05 5.415e-03 1.224e-02
+5 ( 0,2) 2.498e-14 6.278e-08 1.481e-07 5.019e-05 5.351e-03 3.057e-02
+6 ( 1,1) 1.381e-13 1.641e-07 6.657e-07 5.285e-05 6.628e-03 1.612e-02
+7 (-1,1) 1.313e-13 1.148e-07 3.718e-06 7.372e-05 7.391e-03 1.764e-02
+8 ( 2,0) 2.392e-13 7.195e-08 1.482e-07 1.530e-04 8.331e-03 3.442e-02
+9 (-2,0) 4.348e-13 5.816e-08 1.294e-07 2.125e-04 4.244e-03 1.203e-02
+10 ( 0,3) 2.633e-13 4.133e-07 1.408e-06 6.733e-04 3.532e-02 1.011e-01
+11 ( 1,2) 1.293e-12 4.411e-07 7.443e-07 7.967e-04 5.432e-02 1.230e-01
+12 (-1,2) 2.032e-12 2.933e-07 5.402e-07 1.243e-03 8.099e-02 2.104e-01
+13 ( 2,1) 2.639e-12 5.727e-07 1.371e-06 1.452e-03 4.912e-02 2.078e-01
+14 (-2,1) 2.640e-12 6.261e-07 1.510e-06 1.524e-03 2.474e-02 6.840e-02
+15 ( 3,0) 4.376e-12 4.827e-07 8.022e-07 2.384e-03 2.179e-02 7.715e-02
+Table 8: The epochs needed for using SNN-EP with different $M$ to solve the Schrodinger equation for Hydrogen atom.
+M 100 200 300 400 500 600 800 1000 1200
+Epoch 129 91 167 46 127 81 178 119 246
+Figure 5: The dimension of the reduced subspace and the relative error of eigenvalues of the Schrodinger equation for Hydrogen atom obtained by SNN-EP (10 random $\beta_{i}$) with different $M$.
+5 Conclusions
+In this paper, we have proposed a subspace method based on neural networks for eigenvalue problems.
+This method can in fact be viewed as a method combining the machine learning with the classical methods.
+That is, first use the machine learning method and model order reduction method to get some neural network-based orthogonal basis, then calculate the Galerkin projection of the eigenvalue problem onto the subspace spanned by the orthogonal basis and obtain an approximate solution, just like what the classical methods do.
+Our method can achieve high precision with only a few training epochs.
+In fact, from the numerical experiments, we can see that our method can obtain approximate eigenpairs with the error of eigenvalues being below $10^{-11}$ and the $L^{2}$ norm and $H^{1}$ norm error of eigenfunctions being nearly below $10^{-7}$, but with only tens or hundreds of epochs.
+Appendix A. The impact of training on solution accuracy
+In this section, we provide some numerical examples to illustrate how the parameters $\theta$ used in the training affect the solution accuracy.
+After each gradient descent step, we solve the eigenvalue problem in the subspace spanned by the basis functions $\psi_{1}(\bm{x},\theta),\cdots,\psi_{K}(\bm{x},\theta)$ which are obtained by the newly updated parameters $\theta$ and model order reduction.
+We take the results of 2D Laplace eigenvalue problem on a square domain $0,1]^{2}$ as an example, whose details are introduced in Section 4.
+The number of neurons of the subspace layer $M$ is set to 300, and all other parameters are the same with Section 4.
+Figure 6 shows the solution accuracy at different gradient descent steps.
+We can see that for the smallest few eigenvalues, it is easy to obtain high accuracy approximation without training, while for the next smallest eigenvalues, training can significantly improve the accuracy of the results.
+Figure 6: The effect of trained parameters $\theta$ on the accuracy of solution.
+Furthermore, we see the dimension of the subspace spanned by the basis $\psi_{1}(\bm{x},\theta),$ $\cdots,$ $\psi_{K}(\bm{x},\theta)$ generated by training and model order reduction continues to increase as the training steps increase.
+Consequently, more information is preserved during the dimensionality reduction part, which helps to explain the increase in solution accuracy throughout the training process.
+Appendix B. The necessity of employing dimensionality reduction technique
+In this section, we provide some numerical tests that can show the necessity of employment dimensionality reduction technique.
+In fact, we have applied the version of Algorithm 3 that excludes the dimensionality reduction step to solution of 2D Laplace eigenvalue problem on a square domain $0,1]^{2}$ as introduced in Section 4, aiming to solve the first 15 eigenpairs.
+In this situation, the algebraic eigenvalue problem (21) is extremely ill-conditioned.
+Here, we adopt the solver $scipy.linalg.eig$ in Python, which can deal with the ill-conditioned systems in some sense.
+All other settings are exactly the same as in Section 4.
+Numerical experiments for different $M$ have been carried out and the results are presented in Table 9.
+Table 9: The relative errors of the smallest $15$ eigenvalues obtained by the version of Algorithm 3 that excludes the dimensionality reduction step under different $M$ for the 2D Laplace eigenvalue problem on a square domain $0,1]^{2}$.
 Dimension 100 200 300 400 500 600
@@
 $err_{\lambda,15}$ 7.798e-07 4.797e-01 8.226e-02 9.797e-03 4.927e-02 1.188e-02
-After some analysis and numerical verification, we have found that the main reason for the poor accuracy may lie in the extreme ill-conditioning of the matrices $A$ and $B$ in (10).
+From Table 9 we can see that when $M$ is small, our method can obtain approximate eigenvalues with high accuracy.
+However, with the increase of $M$, the accuracy of eigenvalues is not improved but rapidly deteriorates, contrary to our expectations.
+After some tests, we found that the main reason for the poor accuracy may lie in the extreme ill-conditioning of the matrices $A$ and $B$ in (21).
 From Table 10, where the condition numbers of the matrices $A$ and $B$ corresponding to different $M$ are shown, we can see that as the increase of $M$, the condition numbers of matrices $A$ and $B$ increase a lot.
-From Table 1 and Table 2, we believe that it is the increasingly ill-conditioned matrices $A$ and $B$ that prevent the solver from obtaining solutions with high accuracy.
-Table 2: The condition number $\kappa(A)$ and $\kappa(B)$ obtained by Algorithm 2 with different $M$ for the 2D Laplace eigenvalue problem on a square domain $[0,1]^{2}$.
-Dimension 100 200 300 400 500 600
+By taking a detailed look at Table 9 and Table 10, we believe that it is the increasingly ill-conditioned matrices $A$ and $B$ that prevent the solver from obtaining solutions with high accuracy.
+In fact, a similar phenomenon has also been observed in random feature method [12], and the low-rank structure of random feature matrix has been systematically studied in [3, 8, 41, 55].
+Table 10: The condition number $\kappa(A)$ and $\kappa(B)$ generated by the version of Algorithm 3 that excludes the dimensioality reduction step under different $M$ for the 2D Laplace eigenvalue problem on a square domain $0,1]^{2}$.
+M 100 200 300 400 500 600
 $\kappa(A)$ 7.945e+15 2.383e+21 1.754e+26 8.161e+30 3.256e+33 6.145e+33
 $\kappa(B)$ 2.954e+17 4.261e+23 2.024e+28 6.245e+32 8.485e+33 2.101e+34
-4 Dimensionality reduced subspace method based on neural networks for eigenvalue problems
-In the section above, we first simply apply the subspace method based on neural networks proposed in [49] to the solution of eigenvalue problems, and find that it behaves not so well.
-By some numerical tests, we find that the main reason is that as the dimension of the subspace $M$ increases, the condition number of the matrix $A$ and matrix $B$ become extremely large.
-In this case, problem (10) becomes extremely ill-conditioned, which results that solving (10) with a high precision becomes particularly difficult and almost impossible.
-A similar phenomenon has also been observed in random feature method [11], and the low-rank structure of random feature matrix have been systematically studied in [3, 7, 36, 47].
-Therefore, in this section, we propose some strategy to reduce the condition number of the stiffness matrix and the mass matrix.
-In our opinion, the ill-conditioningness of the matrices $A$ and $B$ is resulted by the fact that the obtained basis functions are not completely linearly independent, leading to a lot of redundant information.
-Our basic idea is to find a way to remove the redundant functions, thereby reducing the dimension of the subspace but without significantly reducing the subspace information.
-Based on this idea, we apply some dimensionality reduction technique to the subspace spanned by the neural network based functions as introduced in Section 3, and then obtain an improved version of subspace method based on neural networks for eigenvalue problems, see Algorithm 3 for the details.
-The only difference between Algorithm 2 and Algorithm 3 lies in that in Algorithm 3, a step to reduce the dimension of the subspace is added after obtaining the trained neural network basis functions.
-This step significantly reduces the condition number of the corresponding stiffness matrix and mass matrix, which can improve the solvability of the resulting algebraic eigenvalue problem and then improve the accuracy of the solution.
-After that, we calculate the Galerkin projection of the eigenvalue problem (5) onto this reduced subspace, and finally obtain an approximate solution.
-Algorithm 3 General framework of the dimensionality reduced subspace method based on neural networks for eigenvalue problems
-  1.
-Initialize the neural networks architecture (4), generate randomly $\theta$, and give $w_{1}$$\cdots$,$w_{k}$, then modify the neural network basis functions according to the boundary conditions and the computational domain.
-  2. Update parameters $\theta$ by minimizing some loss function, and obtain the updated basis functions $\varphi_{1},\cdots,\varphi_{M}$.
-  3. Reduce the dimension of the subspace $\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$ by some dimensionality reduction technique.
-  4.Calculate the Galerkin projection of the eigenvalue problem (5) onto the reduced subspace and obtain an approximation of the solution.
-4.1 Dimensionality reduction technique
-In this subsection, we will introduce how we compress a large basis $\{\varphi_{1},\cdots,\varphi_{M}\}$ which numerically lacks good linear independece to a small basis $\{\psi_{1},\cdots,\psi_{K}\}$ with good linear independence while keeping as more necessary information as possible by some dimensionality reduction technique.
-Here, we choose the proper orthogonality decomposition (POD) to achieve this goal.
-We can describe the problem as follows.
-Denote $V=\text{span}\{\varphi_{1},\cdots,\varphi_{M}\}$, $V_{k}\subset V$ any $K$-dimensional subspace of $V$, and $\Pi_{K}$ the orthogonal projection operator from $V$ to $V_{K}$.
-The problem is to find an optimal $K$-dimensional subspace $V_{K}^{\ast}$ to approximate the original space $V$ good enough.
-There will be two subproblems.
-One is that, for a given the dimension of the subspace $K\leq M$, how to find an optimal $K$-dimensional subspace $V_{K}^{\ast}$ to approximate original space $V$ on some least–squares sense, the other is how to choose suitable dimension of the subspace $K$.
-For the first problem, i.e.
-we seek an optimal orthogonal projection $\Pi_{K}^{\ast}:V\rightarrow V_{K}^{\ast}$ minimizing the projection error of $\bm{\Phi}=(\varphi_{1},\cdots,\varphi_{M})^{T}\in[L^{2}(\Omega)]^{M}$ under the following least–squares distance,
-$\|\bm{\Phi}-\Pi_{K}\bm{\Phi}\|^{2}:=\sum\limits_{i=1}^{M}\big{|}\big{|}\varphi_{i}-\sum\limits_{j=1}^{K}(\varphi_{i},\psi_{j})\psi_{j}\big{|}\big{|}_{L^{2}(\Omega)}^{2}$
-Hence, we need to solve the following minimization problem,
-$\min_{\psi_{1},\cdots,\psi_{K}\in L^{2}(\Omega)}\sum\limits_{i=1}^{M}\big{|}\big{|}\varphi_{i}-\sum\limits_{j=1}^{K}(\varphi_{i},\psi_{j})\psi_{j}\big{|}\big{|}_{L^{2}(\Omega)}^{2}\ \text{s.t.}\ (\psi_{i},\psi_{j})=\delta_{ij},$ (13)
-Denote $U=[u_{1},\cdots,u_{K}]\in\mathbb{R}^{M\times K}$ and $\bm{\Psi}=U^{T}\bm{\Phi}.$ We can rewrite the minimization problem (13) to the minimization problem (14) through following simple derivation,
-$\sum\limits_{i=1}^{M}\big{|}\big{|}\varphi_{i}-\sum\limits_{j=1}^{K}(\varphi_{i},\psi_{j})\psi_{j}\big{|}\big{|}_{L^{2}(\Omega)}^{2}=\sum\limits_{i=1}^{M}\big{|}\big{|}\Phi^{T}e_{i}-\Phi^{T}\sum\limits_{j=1}^{K}\langle e_{i},u_{j}\rangle_{\mathcal{M}}u_{j}\big{|}\big{|}_{L^{2}(\Omega)}^{2}=\sum\limits_{i=1}^{M}\big{|}\big{|}e_{i}-\sum\limits_{j=1}^{K}\langle e_{i},u_{j}\rangle_{\mathcal{M}}u_{j}\big{|}\big{|}_{\mathcal{M}}^{2},$
-where $\mathcal{M}=((\varphi_{i},\varphi_{j}))_{i,j}\in\mathbb{R}^{M\times M}$, $\langle x,y\rangle_{\mathcal{M}}=\langle x,\mathcal{M}y\rangle,\|x\|_{\mathcal{M}}=\sqrt{\langle x,x\rangle_{\mathcal{M}}},\forall x,y\in\mathbb{R}^{M}$ and $e_{1},\cdots,e_{M}$ are the standard basis in $\mathbb{R}^{M}$.
-$\min_{u_{1},\cdots,u_{K}\in\mathbb{R}^{M}}\sum\limits_{i=1}^{M}\big{|}\big{|}e_{i}-\sum\limits_{j=1}^{K}\langle e_{i},u_{j}\rangle_{\mathcal{M}}u_{j}\big{|}\big{|}_{\mathcal{M}}^{2}\ \text{s.t.}\ \langle u_{i},u_{j}\rangle_{\mathcal{M}}=\delta_{ij},$ (14)
-Notice that the problem (14) is nothing but the minimization problem considered in the POD method, which has been systematically studied (see [34, 42]), hence we give the following optimal orthogonal projection $\Pi_{K}^{\ast}$ of the problem (14) directly,
-$\mathcal{M}^{2}u_{j}=\lambda_{j}\mathcal{M}u_{j},j=1,\cdots,K,$ (15)
-$\min_{u_{1},\cdots,u_{K}\in\mathbb{R}^{M}}\sum\limits_{i=1}^{M}\big{|}\big{|}e_{i}-\sum\limits_{j=1}^{K}\langle e_{i},u_{j}\rangle_{\mathcal{M}}u_{j}\big{|}\big{|}_{\mathcal{M}}^{2}=\sum\limits_{j=K+1}^{M}\lambda_{j},$ (16)
-where $\lambda_{1}\geq\cdots\lambda_{M}\geq 0$ are the first $M$ eigenvalues of $\mathcal{M}^{2}$ relative to $\mathcal{M}$, $u_{1},\cdots,u_{M}$ are the corresponding eigenvectors.
-The vectors $u_{1},\cdots,u_{K}$ are called POD modes. In practical implementation, we can obtain them by solving the following symmetric $M\times M$ eigenvalue problem
-$\mathcal{M}\bar{u}_{j}=\lambda_{j}\bar{u}_{j},j=1,\cdots,K,$
-$u_{j}=\frac{1}{\sqrt{\lambda_{j}}}\bar{u}_{j},j=1,\cdots,K.$
-For the second problem, since (16) shows the approximate error of the optimal $K$-dimensional subspace is related to the sum of the last $M-K$ eigenvalues of the matrix $\mathcal{M}$ (or the square root of the eigenvlaues), it’s reasonable to choose a small enough $K$ such that the first $K$ eigenvalues (or the square root of the eigenvlaues) account for a high proportion in the overall, i.e.,
-$K=\mathop{\arg\min}\{I(N):I(N)\geq 1-\gamma\},$
-where
-$I(N)=\frac{\sum\limits_{j=1}^{N}\lambda_{j}}{\sum\limits_{j=1}^{M}\lambda_{j}}$ (17)
-or
-$I(N)=\frac{\sum\limits_{j=1}^{N}\sqrt{\lambda_{j}}}{\sum\limits_{j=1}^{M}\sqrt{\lambda_{j}}}$ (18)
-and $\gamma$ is a given parameter.
-In our experiment, we find that the indicator $I(N)$ defined by (18) works better.
-We also have done a series of tests to see the change of condition numbers of stiffness matrix and mass matrix after dimensionality reduction operation.
-The condition numbers of stiffness matrix and mass matrix after the dimensionality reduction operation with $\gamma=10^{-9},10^{-10}$ and $10^{-11}$ are presented in Table 3.
-Comparing Table 3 with Table 2, we see clearly that the condition number of the matrices obtained before the dimensionality reduction operation increases rapidly with the increase of $M$.
-In contrast, the condition numbers of the matrices obtained after the dimensionality reduction operation can be controlled within an acceptable range and do not increase with the dimension of the subspace.
-Various numerical experiments have been done in Section 5.
-It can be observed that solving the eigenvalue problem in the reduced subspace indeed yields high-precision solutions of the eigenvalue problem (5) with an appropriate $\gamma$.
-Table 3: The condition number $\kappa(\bar{A})$ and $\kappa(\bar{B})$ obtained by Algorithm 3 with different $M$ and dimensionality reduction parameter $\gamma$ for the 2D Laplace eigenvalue problem on a square domain $[0,1]^{2}$.
-Condition number Dimension $\gamma=10^{-11}$ $\gamma=10^{-10}$ $\gamma=10^{-9}$
-$\kappa(\bar{A})$ 100 1.00 1.00 1.00
-200 1.00 1.00 1.00
-300 1.00 1.00 1.00
-400 1.00 1.00 1.00
-500 1.00 1.00 1.00
-600 1.00 1.00 1.00
-$\kappa(\bar{B})$ 100 206.26 206.26 199.82
-200 513.95 463.29 353.87
-300 666.72 540.08 369.34
-400 742.69 542.86 376.37
-500 659.94 468.31 374.71
-600 819.00 589.94 451.74
-4.2 Galerkin projection of the original eigenvalue problem onto the reduced subspace
-Now, we have obtained the reduced subspace $\text{span}\{\psi_{1},\psi_{2},\cdots,\psi_{K}\}$.
-We then calculate the Galerkin projection of the original eigenvalue problem (5) onto this subspace.
-From (3) in Section 2, the matrix eigenvalue problem can be formulated as:
-$\bar{A}c=\lambda\bar{B}c,$ (19)
-where $\bar{A}=(a(\psi_{i},\psi_{j}))_{ij},\bar{B}=(b(\psi_{i},\psi_{j}))_{ij}\in\mathbb{R}^{K\times K},c=(c^{(1)},\cdots,c^{(K)})^{T}\in\mathbb{R}^{K}.$
-The first $k$ eigenpairs $(\lambda_{h,1},c_{1}),\cdots,(\lambda_{h,k},c_{k})$ can be obtained by solving the matrix eigenvalue problem (19) using some eigensolver.
-We then calculate the $w$ through $w_{i}=\tilde{R}c_{i},i=1,\cdots,k$, and finally obtain the approximation of the first $k$ eigenpairs $(k\leq K)$ $(\lambda_{h,1},u_{h,1}),\cdots,$ $(\lambda_{h,k},u_{h,k})$, where $u_{h,i}=\sum\limits_{j=1}^{M}w_{i}^{(j)}\varphi_{j}(x,\theta)$, $i=1,\cdots,k.$
-Additionally, although the mass matrix $B$ in (10) is theoretically positive definite, when the $M$ is large, numerical errors during the generation of the mass matrix can lead to some very small negative eigenvalues.
-The dimensionality reduction operation effectively filters out these negative eigenvalues caused by numerical errors, making the mass matrix positive definite.
-Therefore, the resulted algebraic eigenvalue problem is easier to be solved with high accuracy.
-5 Numerical experiments
-In this section, we will use three typical eigenvalue problems, the Laplace eigenvalue problem, the decoupled harmonic oscillator and the coupled harmonic oscillator to demonstrate the performance of the methods proposed in our paper.
-The algorithms are implemented based on the deep learning framework PyTorch.
-All the numerical experiments are carried on the LSSC-IV platform at the Academy of Mathematics and Systems Science, Chinese Academy of Sciences.
-The following parameter settings are used in our numerical experiments: The neural networks architecture is implemented by a fully connected neural network with 3 hidden layers and a subspace layer with each hidden layer containing 100 neurons and the subspace layer containing $M$ neurons.
-We choose the activation function $\sigma(x)=\sin(x).$ During initialization part, the default initialization strategy is adopted to initialize neural network parameters $\theta$ and the random values generated according to a uniform distribution in the range [-1, 1] are applied to initialize coefficients $w_{1},\cdots,w_{k}$.
-The Adam optimizer with the default learning rate is used for training.
-In all experiments, we adopt the termination criteria introduced in (9) with the parameter $N_{max}=5000$.
-Different $\epsilon$ are chosen for different problems to achieve appropriate training (i.e., fewer epochs while maintaining accuracy).
-In the solving part, to improve the accuracy of the solution, we perform dimensionality reduction, with a given parameter denoted as $\gamma$.
-The solver $scipy.linalg.eigh$ in Python is used for solving the matrix eigenvalue problem (19).
-For reproducibility, all the random seeds used in the numerical experiment are set to 1.
-We refer to the dimensionality reduced SNN for eigenvalue problems as R-SNN and compare our method with another neural-network based method proposed in [45], which can be viewed as a variant of DRM capable of solving multiple eigenpairs, simultaneously, denoted as DRM-M.
-For the fairness of the comparison, we choose the same neural networks architecture, initialization method, and the optimizer.
-R-SNN is trained until termination criterion (9) is satisfied, while DRM-M is trained for 25000 epochs.
-The relative error of the approximate eigenvalues and eigenfunctions are defined as follows:
-$err_{\lambda,l}:=\frac{|\lambda_{l}-\lambda_{h,l}|}{|\lambda_{l}|},err_{L^{2},l}:=\frac{\|u_{l}-u_{h,l}\|_{L^{2}(\Omega)}}{\|u_{l}\|_{L^{2}(\Omega)}},err_{H^{1},l}:=\frac{\|u_{l}-u_{h,l}\|_{H^{1}(\Omega)}}{\|u_{l}\|_{H^{1}(\Omega)}},$
-where $\lambda_{l}$ and $u_{l}$ are reference eigenvalues and eigenfucntions, $l=1,\cdots,k$.
-5.1 Laplace eignevalue problem
-[…]
-$\begin{cases}-\Delta u=\lambda u&\text{in }\Omega,\\ u=0&\text{on }\partial\Omega,\end{cases}$
-where $\Omega=[0,1]^{2}$. The exact eigenvalues are as follows:
-$\lambda_{n_{1},n_{2}}=\pi(n_{1}^{2}+n_{2}^{2}),n_{1},n_{2}=1,2,\cdots,$
-and the corresponding eigenfunctions are
-$u_{n_{1},n_{2}}(x_{1},x_{2})=\sqrt{2}\sin\left(\pi n_{1}x_{1}\right)\sin\left(\pi n_{2}x_{2}\right),n_{1},n_{2}=1,2,\cdots.$
-We solve the first $15$ eigenpairs of the 2D Laplace eigenvalue problem using R-SNN.
-Since $\Omega$ is bounded, we choose the Legendre-Gauss integration scheme with 32 integration points in each dimension.
-For training, we set $\epsilon=10^{-3}$.
-And for the dimensionality reduction part, we set $\gamma=10^{-11}$.
-Figure 1: The dimension of the reduced subspace and the relative errors of the first 15 eigenvalues resulted by different $M$ for the 2D Laplace eigenvalue problem obtained by R-SNN.
-The dimension of the reduced subspace and the relative errors of the first 15 eigenvalues with different initial dimension $M$ obtained by R-SNN are presented in Figure 2.
-It can be observed that when the dimension of the original subspace $M$ is small, the dimension of the reduced subspace increases significantly with the increase of $M$.
-At this stage, the relative errors of the first 15 eigenvalues rapidly decay to below $10^{-12}$.
-When $M$ reaches 300, the increase in the dimension of the reduced subspace slows down significantly and the relative errors of the first 15 eigenvalues remain almost unchanged.
-Therefore, we only show results obtained by setting $M=300$ for both R-SNN and DRM-M.
-Table 4: The relative errors of the first 15 eigenvalues with $M$=300 for the 2D Laplace eigenvalue problem.
-Method $k$ ($n_{1},n_{2}$) $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$
-R-SNN 1 (1,1) 3.960e-15 2.799e-08 3.912e-07
-2 (2,1) 4.996e-14 4.711e-08 4.239e-07
-3 (1,2) 1.526e-13 5.802e-08 5.366e-07
-4 (2,2) 1.449e-13 6.954e-08 5.151e-07
-5 (3,1) 1.523e-13 8.539e-08 5.753e-07
-6 (1,3) 1.892e-13 9.419e-08 6.128e-07
-7 (3,2) 1.256e-13 1.137e-07 6.993e-07
-8 (2,3) 3.792e-13 1.284e-07 7.331e-07
-9 (4,1) 9.622e-14 1.416e-07 7.757e-07
-10 (1,4) 2.531e-13 1.436e-07 7.976e-07
-11 (3,3) 1.923e-13 1.774e-07 9.253e-07
-12 (4,2) 1.518e-13 1.737e-07 8.788e-07
-13 (2,4) 5.876e-13 2.258e-07 1.181e-06
-14 (4,3) 2.919e-13 2.558e-07 1.096e-06
-15 (3,4) 4.750e-13 3.231e-07 1.435e-06
-DRM-M 1 (1,1) 3.868e-07 1.222e-04 7.086e-04
-2 (2,1) 1.325e-07 1.135e-04 4.374e-04
-3 (1,2) 2.186e-07 1.249e-04 6.110e-04
-4 (2,2) 7.583e-08 9.439e-05 3.828e-04
-5 (3,1) 1.126e-07 2.027e-04 5.936e-04
-6 (1,3) 1.608e-07 1.709e-04 4.993e-04
-7 (3,2) 8.341e-08 1.899e-04 4.586e-04
-8 (2,3) 1.163e-07 1.825e-04 4.520e-04
-9 (4,1) 3.046e-07 1.360e-03 1.962e-03
-10 (1,4) 1.193e-06 5.145e-04 9.427e-04
-11 (3,3) 1.106e-07 2.519e-04 4.826e-04
-12 (4,2) 2.862e-07 1.372e-03 1.908e-03
-13 (2,4) 1.039e-06 5.832e-04 9.282e-04
-14 (4,3) 3.565e-07 1.184e-03 1.516e-03
-15 (3,4) 6.469e-07 8.103e-04 1.084e-03
-Table 5: The epochs resulted by different $M$ for the 2D Laplace eigenvalue problem obtained by R-SNN.
-Dimension 50 100 150 200 250 300 350 400 450 500 550 600
-Epoch 177 195 190 342 160 165 171 160 271 168 200 191
-The relative errors of the first 15 eigenpairs obtained using R-SNN and DRM-M with the dimension of original subspace $M=300$ are presented in Table 4, and epochs required for our training process are shown in Table 5.
-It can be seen that the accuracy of DRM-M reaches only $10^{-8}$ for eigenvalues and $10^{-4}$ for eigenfunctions after 25000 epochs, while our method achieves accuracy below $10^{-12}$ for eigenvalues and nearly $10^{-7}$ for eigenfunctions, only with 100-400 epochs.
-5.2 Harmonic oscillator problems
-In this subsection, we will study the following 2D decoupled and coupled harmonic oscillator.
-5.2.1 Two-dimensional decoupled harmonic oscillator
-[…]
-$-\frac{1}{2}\Delta u+Vu=\lambda u\quad\text{in}\quad\Omega,$
-where $\Omega=\mathbb{R}^{2},V(x_{1},x_{2})=\frac{1}{2}(x_{1}^{2}+x_{2}^{2})$. The exact eigenvalues are given as follows:
-$\lambda_{n_{1},n_{2}}=(\frac{1}{2}+n_{1})+(\frac{1}{2}+n_{2}),n_{1},n_{2}=0,1,\cdots,$
-and the corresponding eigenfunctions are
-$u_{n_{1},n_{2}}(x_{1},x_{2})=\mathcal{H}_{n_{1}}(x_{1})e^{-x_{1}^{2}/2}\mathcal{H}_{n_{2}}(x_{2})e^{-x_{2}^{2}/2},n_{1},n_{2}=0,1,\cdots,$
-where $\mathcal{H}_{n}$ is Hermite polynomials [4].
-We solve the first $15$ eigenpairs of 2D decoupled harmonic oscillator using R-SNN.
-Since $\Omega$ is $\mathbb{R}^{2}$, the Hermite-Gauss integration scheme is chosen to compute the unbounded integral, with 99 integration points in each dimension.
-In training part, we take $\epsilon=10^{-2}$.
-In the dimensionality reduction part, we take $\gamma=10^{-10}$.
-The dimension of the reduced subspace and the relative errors of the first 15 eigenvalues under different initial dimension $M$ are presented in Figure 3.
-It can be observed that when the original dimension of the subspace is small, the dimension of the reduced subspace increases significantly with the increase of $M$.
-At this stage, the relative errors of the first 15 eigenvalues rapidly decay to below $10^{-11}$.
-When $M$ reaches 900, the increase in the dimension of the reduced subspace slows down significantly and the relative errors of the first 15 eigenvalues remain almost unchanged.
-Therefore, we only show results obtained by setting $M=900$ for both R-SNN and DRM-M.
-Figure 2: The dimension of the reduced subspace and the relative errors resulted by different $M$ for 2D decoupled harmonic oscillator obtained by R-SNN.
-Table 6: The relative errors of the first 15 eigenvalues with $M$=900 for 2D decoupled harmonic oscillator.
-Method $k$ $(n_{1},n_{2})$ $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$
-R-SNN 1 (0,0) 7.294e-14 3.977e-08 2.373e-07
-2 (0,1) 8.005e-14 6.304e-08 3.012e-07
-3 (1,0) 8.926e-14 6.095e-08 2.963e-07
-4 (0,2) 3.449e-14 1.184e-07 4.473e-07
-5 (1,1) 4.737e-14 9.585e-08 3.648e-07
-6 (2,0) 1.198e-13 9.863e-08 4.040e-07
-7 (0,3) 7.283e-14 3.044e-07 1.035e-06
-8 (1,2) 9.104e-14 1.458e-07 5.599e-07
-9 (2,1) 2.474e-13 1.632e-07 5.801e-07
-10 (3,0) 7.412e-13 2.373e-07 8.030e-07
-11 (1,3) 3.457e-12 1.312e-06 3.803e-06
-12 (2,2) 4.765e-12 9.109e-07 2.550e-06
-13 (3,1) 5.116e-12 1.221e-06 3.586e-06
-14 (0,4) 1.962e-11 1.923e-06 5.044e-06
-15 (4,0) 2.506e-11 1.185e-06 3.308e-06
-DRM-M 1 (0,0) 2.449e-09 2.062e-05 3.913e-05
-2 (0,1) 4.052e-09 3.637e-05 6.912e-05
-3 (1,0) 8.495e-09 5.009e-05 9.015e-05
-4 (0,2) 3.324e-09 4.887e-05 7.840e-05
-5 (1,1) 4.008e-09 5.239e-05 8.106e-05
-6 (2,0) 6.147e-09 4.121e-05 6.859e-05
-7 (0,3) 3.407e-09 7.011e-05 1.035e-04
-8 (1,2) 3.609e-09 6.003e-05 8.251e-05
-9 (2,1) 7.057e-09 6.483e-05 9.069e-05
-10 (3,0) 8.040e-09 7.635e-05 1.054e-04
-11 (1,3) 6.594e-09 1.537e-04 1.734e-04
-12 (2,2) 7.201e-09 1.784e-04 2.085e-04
-13 (3,1) 1.029e-08 1.329e-04 1.632e-04
-14 (0,4) 1.426e-08 1.176e-04 1.424e-04
-15 (4,0) 2.001e-08 1.511e-04 1.772e-04
-Table 7: The epochs resulted by different $M$ for 2D decoupled harmonic oscillator obtained by R-SNN.
-Dimension 100 200 300 400 500 600 700 800 900 1200 1500 1800
-Epoch 20 21 23 23 22 20 20 21 21 21 20 22
-The relative errors of the first 15 eigenpairs obtained using R-SNN and DRM-M with the dimension of the original subspace $M=900$ are presented in Table 6.
-Table 7 provides the number of epochs for various $M$.
-Here, it is evident that our accuracy is significantly superior to the results of DRM-M with far fewer epochs.
-Additionally, results from other studies for the same example are also considered.
-An accuracy of $10^{-2}$ is achieved using machine learning methods based on Monte Carlo integration, as demonstrated in [28].
-Furthermore, our accuracy surpasses that of [45], which employs numerical integration based on TNN to achieve high-precision calculations.
-Compared to [45], which requires 500,000 epochs, our method achieves the desired accuracy with only 20 epochs.
-5.2.2 Two-dimensional coupled harmonic oscillator
-[…]
-$-\frac{1}{2}\Delta u+Vu=\lambda u\quad\text{in}\quad\Omega,$
-where $\Omega=\mathbb{R}^{2},V(x_{1},x_{2})=\frac{1}{2}(a_{11}x_{1}^{2}+2a_{12}x_{1}x_{2}+x_{22}^{2})$, the coefficients $a_{11},a_{12},a_{22}$ are selected from [28] and maintains the same accuracy as [28], i.e.
-$a_{11}=0.8851,a_{12}=-0.1382,a_{22}=1.1933$.
-The coupled harmonic oscillator defined by the above potential energy term can be decoupled into the following diagonal form:
-$V(y_{1},y_{2})=\frac{1}{2}(\mu_{1}y_{1}^{2}+\mu_{2}y_{2}^{2}),$
-where $y_{1}=-0.9339352418x_{1}-0.3574422527x_{2}$ and $y_{2}=0.3574422527x_{1}-0.93393524$ $18x_{2}$, $\mu_{1}=0.8322071257$, $\mu_{2}=1.2461928742$.
-Based on the decoupled form, the exact eigenvalues are given as follows:
-$\lambda_{n_{1},n_{2}}=(\frac{1}{2}+n_{1})\mu_{1}^{1/2}+(\frac{1}{2}+n_{2})\mu_{2}^{1/2},n_{1},n_{2}=0,1,\cdots,$
-and the corresponding eigenfunctions are
-$u_{n_{1},n_{2}}(x_{1},x_{2})=\mathcal{H}_{n_{1}}(\mu_{1}^{1/4}y_{1})e^{-\mu_{1}^{1/2}y_{1}^{2}/2}\mathcal{H}_{n_{2}}(\mu_{2}^{1/4}y_{2})e^{-\mu_{2}^{1/2}y_{2}^{2}/2},n_{1},n_{2}=0,1,\cdots.$
-We solve the first $15$ eigenpairs of 2D coupled harmonic oscillator using R-SNN.
-Since $\Omega$ is $\mathbb{R}^{2}$, we choose the Hermite-Gauss integration scheme, with 99 integration points in each dimension.
-In training part, we take $\epsilon=10^{-2}$.
-In the dimensionality reduction part, we take $\gamma=10^{-10}$.
-Figure 3: The dimension of the reduced subspace and the relative errors resulted by different $M$ for 2D coupled harmonic oscillator obtained by R-SNN.
-The dimension of the reduced subspace and the relative errors of the first 15 eigenvalues under different initial dimensions $M$ are presented in Figure 4.
-A numerical performance similar to decoupled harmonic oscillator can be observed.
-Accordingly, we present specific results for $M=900$ for both R-SNN and DRM-M.
-Table 8: The relative errors of the first 15 eigenvalues with $M$=900 for 2D coupled harmonic oscillator.
-Method $k$ $(n_{1},n_{2})$ $err_{\lambda}$ $err_{L^{2}}$ $err_{H^{1}}$
-R-SNN 1 (0,0) 2.330e-11 4.430e-08 2.520e-07
-2 (1,0) 1.979e-11 6.415e-08 3.183e-07
-3 (0,1) 2.647e-11 7.323e-08 3.517e-07
-4 (2,0) 1.939e-11 2.710e-07 1.011e-06
-5 (1,1) 2.348e-11 1.704e-07 6.191e-07
-6 (0,2) 2.740e-11 1.188e-07 4.584e-07
-7 (3,0) 2.829e-11 1.107e-06 3.355e-06
-8 (2,1) 2.312e-11 4.443e-07 1.350e-06
-9 (1,2) 2.545e-11 2.514e-07 8.112e-07
-10 (0,3) 2.798e-11 1.478e-07 5.192e-07
-11 (4,0) 2.524e-10 6.010e-06 1.645e-05
-12 (3,1) 1.159e-10 4.032e-06 1.036e-05
-13 (2,2) 4.021e-11 1.626e-06 4.390e-06
-14 (1,3) 3.324e-11 1.120e-06 2.951e-06
-15 (0,4) 2.874e-11 3.594e-07 1.063e-06
-DRM-M 1 (0,0) 1.154e-07 1.361e-04 2.723e-04
-2 (1,0) 2.964e-08 9.442e-05 1.736e-04
-3 (0,1) 2.188e-08 9.079e-05 1.600e-04
-4 (2,0) 6.870e-08 1.912e-04 2.966e-04
-5 (1,1) 1.830e-08 9.054e-05 1.414e-04
-6 (0,2) 1.826e-08 1.172e-04 1.754e-04
-7 (3,0) 5.469e-08 1.966e-04 2.674e-04
-8 (2,1) 4.192e-08 2.110e-04 2.704e-04
-9 (1,2) 1.655e-08 1.173e-04 1.616e-04
-10 (0,3) 1.759e-08 1.341e-04 1.743e-04
-11 (4,0) 9.014e-08 3.495e-04 4.336e-04
-12 (3,1) 6.991e-08 3.703e-04 4.364e-04
-13 (2,2) 3.048e-08 2.252e-04 2.719e-04
-14 (1,3) 4.191e-08 2.746e-04 3.394e-04
-15 (0,4) 3.972e-08 2.894e-04 3.600e-04
-Table 9: The epochs resulted by different $M$ for 2D coupled harmonic oscillator obtained by R-SNN.
-Dimension 100 200 300 400 500 600 700 800 900 1200 1500 1800
-Epoch 20 21 22 23 22 20 20 21 20 21 20 23
-The relative errors of the first 15 eigenpairs obtained from R-SNN and DRM-M for solving the 2D coupled harmonic oscillator with dimension of the original subspace $M=900$ are shown in Table 8.
-Similar to the decoupled harmonic oscillator, our method provides solutions with higher accuracy while requiring fewer epochs during training, as shown in Table 9, compared to the DRM-M and other existing results in [28] and [45].
-6 Conclusions
-In this paper, we propose a subspace method based on neural networks for eigenvalue problems with high accuracy.
-SNN can in fact be viewed as a method combining both the machine learning method with the traditional methods, that is, use the machine learning method to get some finite dimensional subspace, then calculate the Galerkin projection of the original eigenvalue problems onto the subspace and obtain a approximate solution, just like what the traditional methods do.
-Through analysis and experimental validation, we find that reducing the condition number of the stiffness matrix and the mass matrix through a dimensionality reduction technique can effectively improve the accuracy of approximate solution.
-Our method achieves high precision with only a few training epochs.
-The eigenvalue errors of our method can be below $10^{-11}$, and the eigenfunction errors are nearly below $10^{-7}$, with the required number of epochs ranging from 20 to 400.
-Acknowledgments
-We would like to acknowledge Professor Aihui Zhou for his helpful discussions that improve the presentation of this paper.
+Hence, these results show the necessity of employing some dimensionality reduction technique in Algorithm 3 to construct neural network-based orthogonal basis functions.
+In fact, we have done a series of tests, which show that this step can improve the condition numbers of the stiffness matrix and the mass matrix greatly.
+The condition numbers of the stiffness matrix and the mass matrix with the dimensionality reduction operation with $\gamma=10^{-9},10^{-10}$ and $10^{-11}$ are presented in Table 11.
+Comparing Table 11 with Table 10, we see clearly that the condition number of the matrices obtained without the dimensionality reduction operation increases rapidly with the increase of $M$.
+In contrast, the condition numbers of the matrices obtained with the dimensionality reduction operation can be controlled within an acceptable range and do not increase with $M$.
+Table 11: The condition number $\kappa(A)$ and $\kappa(B)$ obtained by Algorithm 3 with different $M$ and dimensionality reduction parameter $\gamma$ for the 2D Laplace eigenvalue problem on a square domain $0,1]^{2}$.
+$M$ $\kappa(A)$ $\kappa(B)$
+$\gamma=10^{-11}$ $\gamma=10^{-10}$ $\gamma=10^{-9}$ $\gamma=10^{-11}$ $\gamma=10^{-10}$ $\gamma=10^{-9}$
+100 206.26 206.26 199.82 1.00 1.00 1.00
+200 513.95 463.29 353.87 1.00 1.00 1.00
+300 666.72 540.08 369.34 1.00 1.00 1.00
+400 742.69 542.86 376.37 1.00 1.00 1.00
+500 659.94 468.31 374.71 1.00 1.00 1.00
+600 819.00 589.94 451.74 1.00 1.00 1.00